### v1

# Small time asymptotics of spectral heat content of isotropic processes

Abstract.
We provide a general approach for proving small time asymptotic of spectral heat content for any translation invariant isotropic process satisfying negligible tail probability condition. As a consequence, we recover several existing results in the context of Lévy processes and Gaussian processes, and provide spectral heat content asymptotic for a class of time-changed Brownian motions.
Key words and phrases:
spectral heat content; isotropic processes; asymptotic behavior
2020 Mathematics Subject Classification:
60J45; 60G51; 60G15
1. Introduction and main result
Let $X=(X_{t})_{t\geqslant 0}$ be a stochastic process on $\mathbb{R}^{d}$ with càdlàg paths and $\mathbb{P}_{x}$ denotes the probability measure when $X_{0}=x$. For any bounded open set $\Omega\subset\mathbb{R}^{d}$, the spectral heat content of $X$ in the domain $\Omega$ is defined as
$Q^{X}_{\Omega}(t)=\int_{\Omega}\mathbb{P}_{x}(\tau^{X}_{\Omega}>t)dx,$
where $\tau^{X}_{\Omega}=\inf\{t\geqslant 0:X_{t}\notin\Omega\}$. Since $X$ has càdlàg paths, it follows that
$\lim_{t\to 0}Q^{X}_{\Omega}(t)=\mathrm{Vol}(\Omega),$
where $\mathrm{Vol}(\Omega)$ denotes the volume of $\Omega$. In this article, we consider translation invariant isotropic processes $X=(X_{t})_{t\geqslant 0}$ on $\mathbb{R}^{d}$, that is, for any $x\in\mathbb{R}^{d}$, and $A\in\mathcal{B}(\mathbb{R}^{d})$,
$\mathbb{P}_{x}(X_{t}\in A)$ $=\mathbb{P}_{0}(X_{t}+x\in A),$
$\mathbb{P}_{0}(X_{t}\in A)$ $=\mathbb{P}_{0}(UX_{t}\in A)$
for all $d\times d$ orthogonal matrices $U$. Writing $X_{t}=(X^{(1)}_{t},\ldots,X^{(d)}_{t})$, for any $t>0$, we define
(1.1) $\mu(t)$ $=\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\wedge 1\right].$
From right continuity of the sample paths we have
$\lim_{t\to 0}\mu(t)=0.$
Throughout the article, we assume the negligible tail probability condition.
Assumption 1.1.
For any $\varepsilon>0$,
$\lim_{t\to 0}\frac{1}{\mu(t)}\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>\varepsilon\right)=0.$
The goal of this article is to prove the following asymptotic of the spectral heat content of $X$.
Theorem 1.1.
Let $\Omega$ be bounded open subset of $\mathbb{R}^{d}$ with $C^{1,1}$-boundary, and $X=(X_{t})_{t\geqslant 0}$ be a translation invariant isotropic stochastic process satisfying Assumption 1.1. Then,
(1.2) $\lim_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}=\mathrm{Per}(\Omega),$
where $\mathrm{Per}(\Omega)$ denotes the perimeter of $\Omega$.
Remark 1.2.
The $C^{1,1}$ regularity condition on the boundary of $\Omega$ inherits from the differentiability of the signed distance function from the boundary, see Remark 6.6 for details.
Corollary 1.3.
Let $\Omega$ be a bounded, open set with $C^{1,1}$-boundary and $X=(X_{t})_{t\geqslant 0}$ be a translation invariant isotropic process such that
$\mathbb{E}_{0}\left[\left(\sup_{0\leqslant s\leqslant T}X^{(1)}_{s}\right)^{\alpha}\right]<\infty$
for some $T>0$ and $\alpha>1$, and
$\lim_{t\to 0}\frac{\mathbb{E}_{0}\left[\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\right)^{\alpha}\right]}{\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\right]}=0.$
Then,
$\lim_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{m(t)}=\mathrm{Per}(\Omega),$
where $m(t)=\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\right$].
When $X=B$, the standard Brownian motion on $\mathbb{R}^{d}$, the spectral heat content can be realized in terms of the Dirichlet heat equation. More precisely, if $u(t,x)$ solves the PDE
$\frac{\partial u}{\partial t}u(x,t)$ $=\frac{1}{2}\Delta u(x,t)\quad\mbox{in $\Omega\times(0,\infty)$}$
$u(x,t)$ $=0\quad\mbox{for $(x,t)\in\partial\Omega\times(0,\infty)$}$
$u(x,0)$ $=1\quad\mbox{for $x\in\Omega$},$
then the spectral heat content on $\Omega$ for $B$ is given by
$Q^{B}_{\Omega}(t)=\int_{\Omega}u(x,t)dx.$
Small time asymptotic of $Q^{B}_{\Omega}(t)$ was first proved by Berg, see [12, Theorem 6.2], for domains with $R$-smooth boundary, and subsequently, a second order expansion of $Q^{B}_{\Omega}(t)$ near zero was proved in [14, Theorem 1.2] for domains with $C^{3}$-boundary. We also refer to [13] for small time asymptotic of $Q^{B}_{\Omega}$ when $B$ is the Brownian motion on a Riemannian manifold and $\Omega$ has smooth boundary. The spectral heat content asymptotic for symmetric $\alpha$-stable was obtained in [9], and it was later generalized to isotropic Lévy processes in [4], and to symmetric Lévy processes in a recent article [6]. We also refer to [5] for the spectral heat content asymptotic of Gaussian processes in $\mathbb{R}^{d}$.
In the present article, we provide a general approach for spectral heat content asymptotic that works for any translation invariant isotropic process satisfying Assumption 1.1. This assumption holds true for any isotropic Lévy process with regularly varying Lévy-Khintchine exponent, and Gaussian processes. Moreover, in Section 2.3, we provide examples of isotropic processes which are neither Markov nor Gaussian and Theorem 1.1 holds for such processes.
Our method also extends to bounded domains with boundary being level sets of a $C^{1,\kappa}$ functions for $0<\kappa<1$. In all the aforementioned references [12, 13, 14, 9, 4, 5], the assumption of $C^{1,1}$-regularity is crucial as the proofs rely heavily on the uniform interior and exterior ball condition of the domain. In fact, from [7, Corollary 2], the uniform interior and exterior ball condition is equivalent to the $C^{1,1}$ smoothness of the boundary. For $C^{1,\kappa}$-boundary with $0<\kappa<1$, the uniform ball condition does not hold. Theorem 6.1 provides upper and lower bound of the spectral heat content for domains with boundaries given by level sets of $C^{1,\kappa}$ functions with $0<\kappa<1$. The question of validity of (1.2) for such domains still remains open.
We propose a novel approach by considering a functional version of spectral heat content. For any nonnegative integrable function $f:\mathbb{R}^{d}\to\mathbb{R}$ we define
(1.3) $Q^{X}_{f}(t)=\int_{\mathbb{R}^{d}}\mathbb{E}_{x}\left[\inf_{0\leqslant s\leqslant t}f(X_{s})\right]dx,$
In particular, when $f=\mathbbm{1}_{\Omega}$, $Q^{X}_{f}(t)=Q^{X}_{\Omega}(t)$ for all $t\geqslant 0$. First, we obtain asymptotic formula for $Q^{X}_{f}$ when $f$ is a compactly supported smooth function, and the lower bound of the limit in (1.2) is proved using an approximation argument. It turns out that the lower bound of (1.2) holds for any bounded, open subset with finite perimeter. The main idea behind proving the upper bound of the limit in (1.2) stems from the observation that any bounded domain with $C^{1,1}$-boundary can be expressed as $\phi^{-1}(0,\infty)$ for some bounded $C^{1,1}$ function $\phi$ with $|\nabla\phi|$ being constant on the boundary.
The rest of the paper is organized as follows: in Section 2 we provide examples including isotropic Lévy processes, Gaussian processes, and a class of time-changed Brownian motions for which spectral heat content asymptotic follow directly from our main result. Section 3 provides some preliminaries about functions of bounded variation and regularity of boundary. In Section 4, we prove the asymptotic of the functional of the spectral heat content defined in (1.3). Finally, Theorem 1.1 is proved in Section 6.
2. Examples
2.1. Isotropic Lévy processes
We consider the class of Lévy processes studied in [4]. Let $X=(X_{t})_{t\geqslant 0}$ be an isotropic Lévy process with Lévy-Khintchine exponent $\Psi(\xi)=\psi(|\xi|)$ for any $\xi\in\mathbb{R}^{d}$, and assume that $\psi:[0,\infty)\to[0,\infty)$ is regularly varying with exponent $\beta\in[1,2]$. Then, by [10, Equation (3.2)], for any $\varepsilon>0$,
$\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>\varepsilon\right)\leqslant c_{\varepsilon}t$
for some constant $c_{\varepsilon}>0$. On the other hand, using the bounds in [6, Equation (3.8)], we have $\lim_{t\to 0}\frac{t}{\mu(t)}=0$, which shows that Assumption 1.1 holds. Therefore, Theorem 1.1 holds for $X$, and we recover [6, Theorem 2.1(1)] for isotropic Lévy processes. In particular, when $X_{t}$ is the symmetric $\alpha$-stable process, as $t\to 0$ we have
$\mu(t)\sim\begin{cases}t^{1/\alpha}\mathbb{E}[\sup_{0\leqslant s\leqslant 1}X^{(1)}_{s}]&\mbox{if $1<\alpha\leqslant 2$}\\ \frac{1}{\pi}t\log(1/t)&\mbox{if $\alpha=1$}.\end{cases}$
For $\alpha\in(1,2]$, the asymptotic behavior of $\mu(t)$ above follows directly from the self-similarity of $\alpha$-stable process, and for $\alpha=1$, we refer to [2, Proposition 4.3(i)]. This recovers [9, Theorem 1.1] when $\alpha\geqslant 1$.
2.2. Gaussian processes
Consider a translation invariant Gaussian process $X=(X_{t})_{t\geqslant 0}$ on $\mathbb{R}^{d}$ with independent and identically distributed components and assume that $X$ has càdlàg sample paths. For any $t>0$, let
$m(t)=\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\right],\quad\sigma^{2}(t)=\sup_{0\leqslant s\leqslant t}\mathrm{Var}\left[(X^{(1)}_{s})^{2}\right].$
It is known that $m(t)<\infty$ for all $t\geqslant 0$. By Borell-TIS inequality, see [3, Theorem 2.1.1], for any $x>0$
(2.1) $\mathbb{P}_{0}\left(\left|\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}-m(t)\right|>x\right)$ $\leqslant 2e^{-x^{2}/2\sigma^{2}(t)},$
and hence
$\mathrm{Var}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\right]\leqslant\int_{0}^{\infty}4xe^{-x^{2}/2\sigma^{2}(t)}=4\sigma^{2}(t).$
Therefore, $\mathbb{E}_{0}\left[\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\right)^{2}\right]\leqslant m(t)^{2}+4\sigma^{2}(t)$. From [4, Equation (29)] it is known that $\lim_{t\to 0}{\sigma^{2}(t)}/{m(t)}=0$. Applying Corollary 1.3, we obtain the following result.
Proposition 2.1.
If $\Omega$ is a bounded open set with $C^{1,1}$-boundary then,
(2.2) $\lim_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{m(t)}=\mathrm{Per}(\Omega).$
Remark $2.$
In [4, Theorem 4.4] under the condition that the rescaled process $Y^{(t)}_{u}=\mu^{-1}_{t}X^{(1)}_{tu}$ converges weakly to a Gaussian process $(Y_{u})_{u\geqslant 0}$ as $t\downarrow 0$ in Skorokhod $J_{1}$ topology, it is proved that
$\lim_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{m(t)}=\mathbb{E}\left[\sup_{0\leqslant u\leqslant 1}Y_{u}\right]\mathrm{Per}(\Omega),$
with $\mathbb{E}[\sup_{0\leqslant u\leqslant 1}Y_{u}]\leqslant 1$. Proposition 2.1 strengthens [4, Theorem 4.4].
2.3. Time changed Brownian motion
Consider the $d$-dimensional standard Brownian motion $B=(B_{t})_{t\geqslant 0}$ and a nonnegative, increasing continuous process $U=(U_{t})_{t\geqslant 0}$ which is independent of $B$ and $U_{0}=0$. Define $X_{t}=B_{U_{t}}$. Then, $X=(X_{t})_{t\geqslant 0}$ is translation invariant and isotropic. Let us assume that for some $\alpha>1/2$,
(2.3) $\mathbb{E}[U^{\alpha}_{t}]<\infty\quad\forall t\geqslant 0,\quad\lim_{t\to 0}\frac{\mathbb{E}[U^{\alpha}_{t}]}{\mathbb{E}[\sqrt{U_{t}}]}=0.$
Then we have the following result.
Proposition 2.3.
For any bounded open set $\Omega$ with $C^{1,1}$-boundary we have
$\lim_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mathbb{E}[\sqrt{U_{t}}]}=\sqrt{\frac{2}{\pi}}\mathrm{Per}(\Omega).$
Proof.
Since $U_{t}$ is increasing and continuous, for any $t\geqslant 0$ we have
$\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}B^{(1)}_{U_{t}}\right]=\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant U_{t}}B^{(1)}_{s}\right]=\int_{0}^{\infty}\mathbb{E}_{0}\left[\sup_{0\leqslant u\leqslant s}B^{(1)}_{u}\right]\mathbb{P}(U_{t}\in ds).$
Recalling that for any $p>0$,
$\mathbb{E}_{0}\left[\left(\sup_{0\leqslant s\leqslant t}B^{(1)}_{s}\right)^{p}\right]=\frac{2^{p/2}t^{p/2}\Gamma(\frac{p+1}{2})}{\sqrt{\pi}}$
we get
(2.4) $\mathbb{E}_{0}\left[\left(\sup_{0\leqslant s\leqslant t}B^{(1)}_{U_{s}}\right)^{p}\right]=\frac{2^{p/2}\Gamma(\frac{p+1}{2})}{\sqrt{\pi}}\mathbb{E}[U^{p/2}_{t}].$
Therefore, (2.3) implies that the conditions of Corollary 1.3 hold. This concludes the proof. ∎
As an example of $U_{t}$, let us consider an $n$-dimensional standard Brownian motion $W=(W_{t})_{t\geqslant 0}$ independent of $B$ and define the integrated squared radial process
$R_{t}=\int_{0}^{t}|W_{s}|^{2}ds.$
Then, the one-dimensional marginal distributions of $X_{t}=B_{R_{t}}$ coincide with the Lévy stochastic area, see [11, Equation (3.5)]. Also, $X$ is neither Markov nor Gaussian. For any $t\geqslant 0$,
$\mathbb{E}[R_{t}]=\int_{0}^{t}nsds=\frac{nt^{2}}{2},$
and by Cauchy-Schwarz inequality $\mathbb{E}[\sqrt{R_{t}}]\leqslant\sqrt{\mathbb{E}[R_{t}]}<\infty$. On the other hand, using Cauchy-Schwarz inequality, we have
$\sqrt{\frac{R_{t}}{t}}\geqslant\frac{1}{t}\int_{0}^{t}|W_{s}|ds,$
which shows that
$\mathbb{E}[\sqrt{R_{t}}]\geqslant\frac{1}{\sqrt{t}}\mathbb{E}\left[\int_{0}^{t}|W_{s}|ds\right]\geqslant ct$
for some constant $c>0$. Therefore, $\lim_{t\to 0}\mathbb{E}[R_{t}]/\mathbb{E}[\sqrt{R_{t}}]=0$, and hence the conditions in (2.3) are satisfied by $R_{t}$, which implies the following corollary.
Corollary 2.4.
Let the processes $B,R$ be defined as above and $X_{t}=B_{R_{t}}$. Then, for any bounded, open set $\Omega$ with $C^{1,1}$-boundary
$\lim_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mathbb{E}[\sqrt{R_{t}}]}=\sqrt{\frac{2}{\pi}}\mathrm{Per}(\Omega).$
3. Preliminaries: Perimeter and regularity of the boundary
A function $f\in L^{1}(\mathbb{R}^{d})$ has bounded variation if
$\|Df\|:=\sup\left\{\int_{\mathbb{R}^{d}}f\sum_{i=1}^{d}D_{i}\varphi_{i}:\varphi_{i}\in C^{\infty}_{c}(\mathbb{R}^{d}),\sum_{i=1}^{d}\varphi^{2}_{i}\leqslant 1\right\}<\infty.$
In this case, $\|Df\|$ is called the variation of $f$, and the space of all functions with bounded variation is denoted by $\mathrm{BV}(\mathbb{R}^{d})$. A bounded open set $\Omega\subset\mathbb{R}^{d}$ is said to have finite perimeter if $\mathbbm{1}_{\Omega}\in\mathrm{BV}(\mathbb{R}^{d})$, and we write
$\mathrm{Per}(\Omega)=\|D\mathbbm{1}_{\Omega}\|.$
Next, for $\kappa>0$, a function $f:\mathbb{R}^{d}\to\mathbb{R}$ is $C^{1,\kappa}$ if $f$ is $C^{1}$ and $\nabla f$ is uniformly Hölder continuous with exponent $\kappa$, that is, there exists $L>0$ such that
$|\nabla f(x)-\nabla f(y)|\leqslant L|x-y|^{\kappa}$
for all $x,y\in\mathbb{R}^{d}$. The domain $\Omega\subset\mathbb{R}^{d}$ has $C^{1,\kappa}$-boundary if for every $x\in\partial\Omega$, there exists a neighborhood $V_{x}$ and a $C^{1,\kappa}$ function $\gamma:\mathbb{R}^{d-1}\to\mathbb{R}$ such that
$\Omega\cap V_{x}=\{(y_{1},\ldots,y_{d}):y_{d}>\gamma(y_{1},\ldots,y_{d-1})\}\cap V_{x}.$
If $\Omega$ has $C^{1}$-boundary, from Stoke’s theorem it follows that $\mathrm{Per}(\Omega)=\mathcal{H}^{d-1}(\partial\Omega)$, the (d1)(d-1)-dimensional Hausdorff measure of $\partial\Omega$.
4. The minimum functional
Consider a nonnegative measurable function $f$ on $\mathbb{R}^{d}$. For any $u\geqslant 0$, let us define
$E^{f}_{u}=\{x:f(x)>u\}.$
We consider the following functional version of the spectral heat content: for any $f$ as above, we define
$Q^{X}_{f}(t)=\int_{0}^{\infty}Q^{X}_{E^{f}_{u}}(t)du=\int_{0}^{\infty}\int_{\mathbb{R}^{d}}\mathbb{P}_{x}(\tau^{X}_{E^{f}_{u}}>t)dxdu.$
Lemma 4.1.
For any nonnegative measurable function $f$ on $\mathbb{R}^{d}$ one has
$Q^{X}_{f}(t)=\int_{\mathbb{R}^{d}}\mathbb{E}_{x}\left[\inf_{0\leqslant s\leqslant t}f(X_{s})\right]dx.$
In particular, when $f=\mathbbm{1}_{\Omega}$, $Q^{X}_{f}(t)=Q^{X}_{\Omega}(t)$.
Proof.
We note that for any fixed $u\geqslant 0$,
$\left\{u<\inf_{0\leqslant s\leqslant t}f(X_{s})\right\}\subseteq\{\tau_{E^{f}_{u}}>t\}=\left\{\inf_{s\geqslant 0}\{s:f(X_{s})\leqslant u\}>t\right\}\subseteq\left\{u\leqslant\inf_{0\leqslant s\leqslant t}f(X_{s})\right\}$
Applying Fubini’s theorem,
$Q^{X}_{f}(t)=\int_{\mathbb{R}^{d}}\int_{0}^{\infty}\mathbb{P}_{x}\left(\inf_{0\leqslant s\leqslant t}f(X_{s})\geqslant u\right)dudx=\int_{\mathbb{R}^{d}}\mathbb{E}_{x}\left[\inf_{0\leqslant s\leqslant t}f(X_{s})\right]dx.$
This completes the proof of the lemma. ∎
In what follows, we will study the asymptotic behavior of the quantity
$\mathcal{R}^{X}_{f}(t):=\int_{\mathbb{R}^{d}}f(x)dx-Q^{X}_{f}(t).$
as $t\to 0$. Trivially, when $f=\mathbbm{1}_{\Omega}$, $\mathcal{R}^{X}_{f}(t)=\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)$.
Theorem 4.2.
Let $X=(X_{t})_{t\geqslant 0}$ be a translation invariant, isotropic stochastic process on $\mathbb{R}^{d}$ satisfying Assumption 1.1. Then for any nonnegative $f\in C^{1,\kappa}_{c}(\mathbb{R}^{d})$,
$\lim_{t\to 0}\frac{\mathcal{R}^{X}_{f}(t)}{\mu(t)}=\int_{\mathbb{R}^{d}}|\nabla f(x)|dx,$
where $\mu(t)$ is defined in (1.1). In particular, the above limit holds for all compactly supported smooth functions.
We require the following lemmas to prove the above theorem.
Lemma 4.3.
For any $\varepsilon>0$,
(4.1) $\lim_{t\to 0}\frac{1}{\mu(t)}\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\wedge\varepsilon\right]=1$
(4.2) $\lim_{t\to 0}\frac{1}{\mu(t)}\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\mathbbm{1}\left\{\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\leqslant\varepsilon\right\}\right]=1$
(4.3) $\lim_{t\to 0}\frac{1}{\mu(t)}\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\mathbbm{1}\left\{\sup_{0\leqslant s\leqslant t}|X_{s}|\leqslant\varepsilon\right\}\right]=1$
(4.4) $\limsup_{t\to 0}\frac{1}{\mu(t)}\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}|X_{s}|\mathbbm{1}\left\{\sup_{0\leqslant s\leqslant t}|X_{s}|\leqslant\varepsilon\right\}\right]\leqslant 2d$
Proof.
For any $\varepsilon>0$ we note that
$\left|\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\wedge\varepsilon\right]-\mu(t)\right|\leqslant\max\{\varepsilon,1\}\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>\varepsilon\wedge 1\right).$
Therefore, (4.1) follows as $\lim_{t\to 0}\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>\varepsilon\wedge 1\right)/\mu(t)=0$. Next, for any $t>0$, we have
$0$ $\leqslant\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\wedge\varepsilon\right]-\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\mathbbm{1}\left\{\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\leqslant\varepsilon\right\}\right]$
$\leqslant\varepsilon\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>\varepsilon\right),$
and hence, (4.2) follows from (4.1) and Assumption 1.1. To prove (4.3), we observe that
$0$ $\leqslant\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\leqslant\varepsilon\}\right]-\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}|X_{s}|\leqslant\varepsilon\}\right]$
$=\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\leqslant\varepsilon,\sup_{0\leqslant s\leqslant t}|X_{s}|>\varepsilon\}\right]$
$\leqslant\varepsilon\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}|X_{s}|>\varepsilon\right)\leqslant 2d\varepsilon\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>\varepsilon/\sqrt{d}\right),$
where the last inequality follows from the isotropy of $X_{t}$. Again by Assumption 1.1 and (4.2), we conclude that
(4.5) $\ \ \ \ \lim_{t\to 0}\frac{1}{\mu(t)}\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}|X_{s}|\leqslant\varepsilon\}\right]$
$=\lim_{t\to 0}\frac{1}{\mu(t)}\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\leqslant\varepsilon\}\right]=1.$
Finally, using the symmetry of the random variables $X^{(i)}_{t}$ for $i=1,\ldots,d$, we note that
$\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}|X^{(i)}_{s}|\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}|X^{(i)}_{s}|\leqslant\varepsilon\}\right]\leqslant\int_{0}^{\varepsilon}\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}|X^{(i)}_{s}|>u\right)du$
$\leqslant 2\int_{0}^{\varepsilon}\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}X^{(i)}_{s}>u\right)=2\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(i)}_{s}\wedge\varepsilon\right].$
[…]
$\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}|X_{s}|\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}|X_{s}|\leqslant\varepsilon\}\right]\leqslant\sum_{i=1}^{d}\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}|X^{(i)}_{s}|\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}|X^{(i)}_{s}|\leqslant\varepsilon\}\right].$
Therefore, (4.4) follows from (4.5) and (4.3). ∎
Lemma 4.4.
Suppose that $f\in C^{1,\kappa}(\mathbb{R}^{d})$ with $|\nabla f(x)-\nabla f(y)|\leqslant L|x-y|^{\kappa}$ for all $x,y\in\mathbb{R}^{d}$, and $t>0$. Then,
$\left|\inf_{0\leqslant s\leqslant t}f(X_{s})-f(x)-\inf_{0\leqslant s\leqslant t}\langle X_{s}-x,\nabla f(x)\rangle\right|\leqslant\frac{L}{1+\kappa}\sup_{0\leqslant s\leqslant t}|X_{s}-x|^{1+\kappa}$
Proof.
Since $f\in C^{1}(\mathbb{R}^{d})$, using Taylor expansion around $x$, we obtain,
(4.6) $f(X_{s})=f(x)+\langle X_{s}-x,\nabla f(x)\rangle+\omega_{f}(x,X_{s})$
where $\omega_{f}(x,x^{\prime})=\int_{0}^{1}\langle(x^{\prime}-x),(\nabla f(x+u(x^{\prime}-x))-\nabla f(x))\rangle du$. Using the $\kappa$-Hölder continuity of $\nabla f$, it follows that
(4.7) $|\omega_{f}(x,x^{\prime})|\leqslant\frac{L}{1+\kappa}|x-x^{\prime}|^{1+\kappa}.$
This completes the proof of the lemma. ∎
Proof of Theorem 4.2.
Let us write $S_{f}=\mathrm{supp}(f)$. Then we observe that
$f(x)-\mathbb{E}_{x}\left[\inf_{0\leqslant s\leqslant t}f(X_{s})\right]=0$
whenever $x\notin S_{f}$. As a result, using (4.6), $\mathcal{R}^{X}_{f}(t)$ reduces to
(4.8) $\mathcal{R}^{X}_{f}(t)$ $=\int_{S_{f}}\mathbb{E}_{x}\left[-\inf_{0\leqslant s\leqslant t}\left(\langle X_{s}-x,\nabla f(x)\rangle+\omega_{f}(X_{s},x)\right)\mathbbm{1}\{X_{s}\in S_{f},0\leqslant s\leqslant t\}\right]dx$
Assuming that $S_{f}\subset B_{R}(0)$ for some $R>0$, (4.8) implies
(4.9) $\left|\mathcal{R}^{X}_{f}(t)-\int_{S_{f}}\mathbb{E}_{x}\left[\sup_{0\leqslant s\leqslant t}-\langle X_{s}-x,\nabla f(x)\rangle\mathbbm{1}\{X_{s}\in S_{f},0\leqslant s\leqslant t\}\right]dx\right|$
$\leqslant\int_{S_{f}}\mathbb{E}_{x}\left[\sup_{0\leqslant s\leqslant t}|\omega_{f}(X_{s},x)|\mathbbm{1}\left\{\sup_{0\leqslant s\leqslant t}|X_{s}-x|\leqslant 2R\right\}\right]dx$
$\leqslant\frac{L}{1+\kappa}\int_{S_{f}}\mathbb{E}_{x}\left[\sup_{0\leqslant s\leqslant t}|X_{s}-x|^{1+\kappa}\mathbbm{1}\left\{\sup_{0\leqslant s\leqslant t}|X_{s}-x|\leqslant 2R\right\}\right]dx,$
where we used the bound in (4.7). For $\varepsilon>0$, the translation invariance and isotropy of $X$ leads to
$\ \ \ \ \mathbb{E}_{x}\left[\sup_{0\leqslant s\leqslant t}|X_{s}-x|^{1+\kappa}\mathbbm{1}\left\{\sup_{0\leqslant s\leqslant t}|X_{s}-x|\leqslant 2R\right\}\right]$
$\leqslant\varepsilon^{\kappa}\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}|X_{s}|\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}|X_{s}|\leqslant\varepsilon\}\right]+2^{1+\kappa}R^{1+\kappa}\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}|X_{s}|>\varepsilon\right).$
This shows with Assumption 1.1 and (4.4) that for any $\varepsilon>0$,
$\limsup_{t\to 0}\frac{1}{\mu(t)}\int_{S_{f}}\mathbb{E}_{x}\left[\sup_{0\leqslant s\leqslant t}|X_{s}-x|^{1+\kappa}\mathbbm{1}\left\{\sup_{0\leqslant s\leqslant t}|X_{s}-x|\leqslant 2R\right\}\right]dx\leqslant 2d\varepsilon^{\kappa}\mathrm{Vol}(S_{f}).$
Letting $\varepsilon\to 0$, the above inequality combined with (4.9) yields
(4.10) $\lim_{t\to 0}\frac{1}{\mu(t)}\left|\mathcal{R}^{X}_{f}(t)-\int_{S_{f}}\mathbb{E}_{x}\left[\sup_{0\leqslant s\leqslant t}-\langle X_{s}-x,\nabla f(x)\rangle\mathbbm{1}\{X_{s}\in S_{f},0\leqslant s\leqslant t\}\right]dx\right|=0.$
Let us denote
$\delta(x)=\inf\{|x-y|:y\in\partial S_{f}\}.$
Then, any $x\in S_{f}$ and $t\geqslant 0$ we have
(4.11) $\ \ \ \sup_{0\leqslant s\leqslant t}-\langle X_{s}-x,\nabla f(x)\rangle\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}|X_{s}-x|\leqslant\delta(x)\}$
$\leqslant\sup_{0\leqslant s\leqslant t}-\langle X_{s}-x,\nabla f(x)\rangle\mathbbm{1}\{X_{s}\in S_{f},0\leqslant s\leqslant t\}$
$\leqslant\sup_{0\leqslant s\leqslant t}-\langle X_{s}-x,\nabla f(x)\rangle\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}|X_{s}-x|\leqslant 2R\}$
Recalling the isotropy and translation invariance of $X$, for any $r>0$ we have
$\ \ \ \ \mathbb{E}_{x}\left[\sup_{0\leqslant s\leqslant t}-\langle X_{s}-x,\nabla f(x)\rangle\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}|X_{s}-x|\leqslant r\}\right]$
$=\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}|X_{s}|\leqslant r\}\right]|\nabla f(x)|=:\mu(t,r)|\nabla f(x)|.$
As a result, (4.11) implies the following upper and lower bound:
(4.12) $\mu(t,\delta(x))|\nabla f(x)|$
$\leqslant\mathbb{E}_{x}\left[\sup_{0\leqslant s\leqslant t}-\langle X_{s}-x,\nabla f(x)\rangle\mathbbm{1}\{X_{s},x\in S_{f},0\leqslant s\leqslant t\}\right]$
$\leqslant\mu(t,2R)|\nabla f(x)|.$
Combining (4.10) and (4.12), we obtain
(4.13) $\limsup_{t\to 0}\frac{\mathcal{R}^{X}_{f}(t)}{\mu(t)}$ $\leqslant\limsup_{t\to 0}\frac{\mu(t,2R)}{\mu(t)}\int_{\Omega}|\nabla f(x)|dx$
$=\int_{\mathbb{R}^{d}}|\nabla f(x)|dx,$
where the last equality follows from (4.2) in Lemma 4.3. On the other hand, using (4.10) and (4.12) once again we get
$\liminf_{t\to 0}\frac{\mathcal{R}^{X}_{f}(t)}{\mu(t)}\geqslant\liminf_{t\to 0}\int_{S_{f}}|\nabla f(x)|\frac{\mu(t,\delta(x))}{\mu(t)}dx.$
Since by (4.2) in Lemma 4.3, $\frac{\mu(t,\delta(x))}{\mu(t)}\to 1$ as $t\to 0$, applying Fatou’s lemma we conclude that
(4.14) $\liminf_{t\to 0}\frac{\mathcal{R}^{X}_{f}(t)}{\mu(t)}\geqslant\int_{S_{f}}|\nabla f(x)|dx=\int_{\mathbb{R}^{d}}|\nabla f(x)|dx.$
Combining (4.13) and (4.14), we conclude the proof of the theorem. ∎
5. Proof of the lower bound
Proposition 5.1.
Let $\Omega$ be a bounded, open set in $\mathbb{R}^{d}$ with finite perimeter. Then,
$\liminf_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\geqslant\mathrm{Per}(\Omega).$
Remark 5.2.
When $\Omega$ has $C^{1,1}$-boundary, the above proposition can be proved using a half-space argument similar to the proof of [9, Lemma 3.1]. However, the argument in the aforementioned paper fails when $\partial\Omega$ is not regular.
The above result is proved using approximation argument using mollifiers. Fix a collection of mollifiers $(\rho_{\varepsilon})_{\varepsilon>0}\in C^{\infty}_{c}(\mathbb{R}^{d})$ such that $\int_{\mathbb{R}^{d}}\rho_{\varepsilon}(x)dx=1$ and for any $\varepsilon>0$ and $f\in L^{1}(\mathbb{R}^{d})$, define the convolution
$f_{\varepsilon}=f\ast\rho_{\varepsilon}.$
Lemma 5.3.
For any $f\in L^{1}(\mathbb{R}^{d})$ with $f\geqslant 0$ and $\varepsilon>0$, we have
$\mathcal{R}^{X}_{f_{\varepsilon}}(t)\leqslant\mathcal{R}^{X}_{f}(t)$
for all $t\geqslant 0$, where $f_{\varepsilon}=f\ast\rho_{\varepsilon}$.
Proof.
We note that $\int_{\mathbb{R}^{d}}f_{\varepsilon}(x)dx=\int_{\mathbb{R}^{d}}f(x)dx$ for all $\varepsilon>0$. Also, by translation invariance of $X_{t}$ and Fubini’s theorem we get
$\mathbb{E}_{x}\left[\inf_{0\leqslant s\leqslant t}f_{\varepsilon}(X_{s})\right]$ $=\mathbb{E}_{0}\left[\inf_{0\leqslant s\leqslant t}f_{\varepsilon}(X_{s}+x)\right]$
$=\mathbb{E}_{0}\left[\inf_{0\leqslant s\leqslant t}\int_{\mathbb{R}^{d}}f(X_{s}+x-y)\rho_{\varepsilon}(y)dy\right]$
$\geqslant\mathbb{E}_{0}\left[\int_{\mathbb{R}^{d}}\inf_{0\leqslant s\leqslant t}f(X_{s}+x-y)\rho_{\varepsilon}(y)dy\right]$
$=\int_{\mathbb{R}^{d}}\mathbb{E}_{x-y}\left[\inf_{0\leqslant s\leqslant t}f(X_{s})\right]\rho_{\varepsilon}(y)dy.$
This shows that for all $\varepsilon>0$,
$\int_{\mathbb{R}^{d}}\mathbb{E}_{x}\left[\inf_{0\leqslant s\leqslant t}f_{\varepsilon}(X_{s})\right]dx\geqslant\int_{\mathbb{R}^{d}}\mathbb{E}_{x}\left[\inf_{0\leqslant s\leqslant t}f(X_{s})\right]dx.$
This completes the proof of the lemma. ∎
Proof of Proposition 5.1.
Let us consider $f=\mathbbm{1}_{\Omega}$. By Lemma 5.3 and Theorem 4.2, for any $\varepsilon>0$,
$\liminf_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\geqslant\liminf_{t\to 0}\frac{\mathcal{R}^{X}_{f_{\varepsilon}}(t)}{\mu(t)}=\int_{\mathbb{R}^{d}}|\nabla f_{\varepsilon}(x)|dx.$
As $\lim_{\varepsilon\to 0}\|f_{\varepsilon}-\mathbbm{1}_{\Omega}\|_{L^{1}}=0$, by lower semi-continuity of the variation it follows that
$\liminf_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\geqslant\liminf_{\varepsilon\to 0}\int_{\mathbb{R}^{d}}|\nabla f_{\varepsilon}(x)|\mu(dx)\geqslant\|D\mathbbm{1}_{\Omega}\|=\mathrm{Per}(\Omega).$
This completes the proof of the proposition. ∎
6. Proof of Theorem 1.1
We prove a stronger version of Theorem 1.1.
Theorem 6.1.
Let $\phi:\mathbb{R}^{d}\to\mathbb{R}$ be a $C^{1,\kappa}$ function with $0<\kappa<1$ and assume that
  1. (i)
    $\Omega=\phi^{-1}(0,\infty)$ is pre-compact,
  2. (ii)
    $|\nabla\phi|$ is bounded away from 0 in a neighborhood of $\partial\Omega$.
Then for any translation invariant isotropic process $X=(X_{t})_{t\geqslant 0}$ satisfying Assumption 1.1 we have
$\mathrm{Per}(\Omega)\leqslant\liminf_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\leqslant\limsup_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\leqslant\frac{\sup_{\partial\Omega}|\nabla\phi|}{\inf_{\partial\Omega}|\nabla\phi|}\mathrm{Per}(\Omega).$
The theorem is proved in several intermediate steps. First, we note the following result about continuity of the perimeters of level sets.
Lemma 6.2.
Let $\phi:\mathbb{R}^{d}\to\mathbb{R}$ be a $C^{1}$ function such that
  1. (i)
    $\Omega=\phi^{-1}(0,\infty)$ is pre-compact,
  2. (ii)
    $|\nabla\phi|$ is bounded away from 0 in a neighborhood of $\partial\Omega$.
For any $r>0$, define $\Omega_{r}=\phi^{-1}(r,\infty)$. Then,
$\lim_{r\to 0}\mathcal{H}^{d-1}(\partial\Omega_{r})=\mathcal{H}^{d-1}(\partial\Omega)$
Notation 6.1.
For the rest of the article, we denote
$\Omega^{\prime}_{r}=\{x:0<\phi(x)<r\},\quad\Omega_{r}=\{x:\phi(x)>r\},\quad r>0.$
Then, $\partial\Omega_{r}=\phi^{-1}(r)$.
For a proof of the above lemma, we refer the reader to the Mathoverflow post [1] and answer by Mathoverflow user Kostya_I, where the lemma has been proved for smooth functions $\phi$, but the same argument works for $C^{1}$ functions as the proof only relies on inverse function theorem. For the sake of completeness, we provide the proof below with some minor modifications.
Proof.
We claim that every point $x\in\phi^{-1}(0)$ has a neighborhood $V_{x}$ that is a diffeomorphic image $\varphi_{x}(Q_{r_{x}})$ of a box $Q_{r_{x}}=\{|y_{i}|\leqslant r_{x},\;1\leqslant i\leqslant d\}$, such that $\phi(\varphi(y))=|\nabla\phi(x)|\cdot y_{1}$, and such that $\varphi_{x}(0)=x$ and $D\varphi_{x}(0)$ is a rotation. Indeed, without loss of generality we may assume that $x=0$, and, by rotating, that $\nabla\phi(0)/|\nabla\phi(0)|=e_{1},$ the first basis vector. Consider the function $g:\mathbb{R}^{d}\to\mathbb{R}^{d}$ defined by $g(x)=(\phi(x)/|\nabla\phi(0)|,x_{2},\dots,x_{d})$. Then $Dg(0)$ is the identity, and $\varphi_{x}$ is the inverse $g^{-1}$ provided by the inverse function theorem. Given $\varepsilon>0$, by choosing smaller $r_{x}$ if necessary, we can ensure that the restriction of $\varphi_{x}$ to the leaves $\{y_{1}=r\}$ distorts the (d1)(d-1)-area by no more than $1+\varepsilon$, that is, in $Q_{\delta_{x}}$,
(6.1) $1-\varepsilon<\left(\det_{2\leqslant i,j\leqslant d}\left(\partial_{y_{i}}\varphi\cdot\partial_{y_{j}}\varphi\right)\right)^{\frac{1}{2}}\leqslant 1+\varepsilon.$
By compactness, we can choose a finite cover $V_{i}=V_{x_{i}}$, $i=1,\dots,n$ of $\phi^{-1}(0)$. Then, for $\varepsilon_{0}>0$ small enough, we have $\overline{\Omega}^{\prime}_{\varepsilon_{0}}:=\{x:0\leqslant\phi(x)\leqslant\varepsilon_{0}\}\subset\cup_{i=1}^{n}V_{i}$. Let $V_{1},\dots,V_{n}$ be a finite cover of $\Omega^{\prime}_{\varepsilon_{0}}$, and we can pick a partition of unity $f_{0},\dots,f_{n}$ subordinate to that cover. We have for $0<r<\varepsilon_{0}$,
$\mathcal{H}^{d-1}(\phi^{-1}(r))=\sum_{i=1}^{n}\int_{\phi^{-1}(r)}f_{i}(w)\mathcal{H}^{d-1}(dw).$
Hence, by changing the variable and (6.1),
(6.2) $(1-\varepsilon)\sum_{i=1}^{n}\int_{Q^{r}_{i}}f_{i}\circ\varphi_{i}dy_{2}\ldots dy_{d}\leqslant\mathcal{H}^{d-1}(\phi^{-1}(r))\leqslant(1+\varepsilon)\sum_{i=1}^{n}\int_{Q^{r}_{i}}f_{i}\circ\varphi_{i}dy_{2}\ldots dy_{d},$
where $Q^{r}_{i}=Q_{i}\cap\{y_{1}=r\}$. As $r\to 0$, the terms in the right-hand side tend to
$(1+\varepsilon)\int_{Q^{0}_{i}}f_{i}\circ\varphi_{i}dy_{2}\ldots dy_{d}\leqslant(1+\varepsilon)^{2}\int_{\phi^{-1}(0)}f_{i}(w)\mathcal{H}^{d-1}(dw),$
and similarly for the left hand side of (6.2). This means that
$(1-\varepsilon)^{2}\mathcal{H}^{d-1}(\phi^{-1}(0))$ $\leqslant\liminf_{r\to 0}\mathcal{H}^{d-1}(\phi^{-1}(r)),$
$\limsup_{r\to 0}\mathcal{H}^{d-1}(\phi^{-1}(r))$ $\leqslant(1+\varepsilon)^{2}\mathcal{H}^{d-1}(\phi^{-1}(0)),$
and since $\varepsilon>0$ is arbitrary and $\phi^{-1}(r)=\partial\Omega_{r}$ for any $r\geqslant 0$, the proof the lemma follows. ∎
Lemma 6.3.
For any $\varepsilon>0$,
$\lim_{t\to 0}\frac{1}{\mu(t)}\int_{\Omega_{\varepsilon}}\mathbb{P}_{x}\left(\inf_{0\leqslant s\leqslant t}\phi(X_{s})\leqslant 0\right)dx=0$
Proof.
Invoking Lemma 4.4 we have
$\inf_{0\leqslant s\leqslant t}\phi(X_{s})\geqslant\phi(x)+\inf_{0\leqslant s\leqslant t}\langle X_{s}-x,\nabla\phi(x)\rangle-\frac{L}{1+\kappa}\sup_{0\leqslant s\leqslant t}|X_{s}-x|^{1+\kappa}.$
Therefore,
$\ \ \ \ \mathbb{P}_{x}\left(\inf_{0\leqslant s\leqslant t}\phi(X_{s})\leqslant 0\right)$
$\leqslant\mathbb{P}_{x}\left(\sup_{0\leqslant s\leqslant t}\langle X_{s}-x,\nabla\phi(x)\rangle+\frac{L}{1+\kappa}\sup_{0\leqslant s\leqslant t}|X_{s}-x|^{1+\kappa}>\phi(x)\right)$
$\leqslant\mathbb{P}_{x}\left(\sup_{0\leqslant s\leqslant t}|X_{s}-x|>\frac{\phi(x)}{2M}\right)+\mathbb{P}_{x}\left(\sup_{0\leqslant s\leqslant t}|X_{s}-x|>\left(\frac{\phi(x)}{L^{\prime}}\right)^{1/1+\kappa}\right),$
where $M=\|\nabla\phi\|_{\infty}$ and $L^{\prime}=L/(1+\kappa)$. As a result, for any $x\in\Omega_{\varepsilon}$, Assumption 1.1 implies that
$\ \ \ \ \lim_{t\to 0}\frac{1}{\mu(t)}\int_{\Omega_{\varepsilon}}\mathbb{P}_{x}\left(\inf_{0\leqslant s\leqslant t}\phi(X_{s})\leqslant 0\right)dx$
$\leqslant\lim_{t\to 0}\frac{1}{\mu(t)}\left(\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}|X_{s}|>\frac{\varepsilon}{2M}\right)+\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}|X_{s}|>\left(\frac{\varepsilon}{L^{\prime}}\right)^{1/1+\kappa}\right)\right)\mathrm{Vol}(\Omega)$
$=0.$
This completes the proof of the lemma. ∎
Lemma 6.4.
For any $\eta\in(0,1)$ and $\varepsilon>0$,
$\limsup_{t\to 0}\frac{1}{\mu(t)}\int_{\Omega^{\prime}_{\varepsilon}}\mathbb{P}_{x}\left(\inf_{0\leqslant s\leqslant t}\phi(X_{s})\leqslant 0\right)dx$
$\leqslant\sup_{0\leqslant v\leqslant\varepsilon}\mathcal{H}^{d-1}(\partial\Omega_{v})\frac{M_{\varepsilon}}{\eta m_{\varepsilon}}+2\sup_{0\leqslant v\leqslant\varepsilon}\mathcal{H}^{d-1}(\partial\Omega_{v})\frac{(1-\eta)^{-\frac{\kappa}{\kappa+1}}}{m_{\varepsilon}}\varepsilon^{\frac{\kappa}{\kappa+1}},$
where $M_{\varepsilon}=\sup_{x\in\Omega^{\prime}_{\varepsilon}}|\nabla\phi(x)|$ and $m_{\varepsilon}=\inf_{x\in\Omega^{\prime}_{\varepsilon}}|\nabla\phi(x)|$.
Proof.
Fix any $\eta\in(0,1)$. For any $x\in\Omega^{\prime}_{\varepsilon}$, using Lemma 4.4 again we get
$\ \ \ \ \mathbb{P}_{x}\left(\inf_{0\leqslant s\leqslant t}\phi(X_{s})\leqslant 0\right)$
(6.3) $\leqslant\mathbb{P}_{x}\left(\sup_{0\leqslant s\leqslant t}\langle X_{s}-x,\nabla\phi(x)\rangle>\eta\phi(x)\right)+\mathbb{P}_{x}\left(\sup_{0\leqslant s\leqslant t}|X_{s}-x|^{1+\kappa}>(1-\eta)\phi(x)\right).$
Using translation invariance and isotropy of $X$, and the co-area formula, integral of the first term on $\Omega^{\prime}_{\varepsilon}$ reduces to
(6.4) $\ \ \ \ \int_{\Omega^{\prime}_{\varepsilon}}\mathbb{P}_{x}\left(\sup_{0\leqslant s\leqslant t}\langle X_{s}-x,\nabla\phi(x)\rangle>\eta\phi(x)\right)dx$
$=\int_{0}^{\varepsilon}\int_{\partial\Omega_{v}}\mathbb{P}_{0}\left(|\nabla\phi(x)|\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>\eta v\right)\frac{\mathcal{H}^{d-1}(dx)}{|\nabla\phi(x)|}dv.$
Then, (6.4) yields
$\ \ \ \ \int_{\Omega^{\prime}_{\varepsilon}}\mathbb{P}_{x}\left(\sup_{0\leqslant s\leqslant t}\langle X_{s}-x,\nabla\phi(x)\rangle>\eta\phi(x)\right)dx$
$\leqslant\sup_{0\leqslant v\leqslant\varepsilon}\mathcal{H}^{d-1}(\partial\Omega_{v})\frac{1}{m_{\varepsilon}}\int_{0}^{\varepsilon}\mathbb{P}_{0}\left(M_{\varepsilon}\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>\eta v\right)dv$
$=\sup_{0\leqslant v\leqslant\varepsilon}\mathcal{H}^{d-1}(\partial\Omega_{v})\frac{M_{\varepsilon}}{\eta m_{\varepsilon}}\mathbb{E}_{0}\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\wedge\frac{\eta\varepsilon}{M_{\varepsilon}}\right)$
By (4.1) in Lemma 4.3, we get
(6.5) $\limsup_{t\to 0}\frac{1}{\mu(t)}\int_{\Omega^{\prime}_{\varepsilon}}\mathbb{P}_{x}\left(\sup_{0\leqslant s\leqslant t}\langle X_{s}-x,\nabla\phi(x)\rangle>\eta\phi(x)\right)dx$
$\leqslant\sup_{0\leqslant v\leqslant\varepsilon}\mathcal{H}^{d-1}(\partial\Omega_{v})\frac{M_{\varepsilon}}{\eta m_{\varepsilon}}$
To estimate the integral of the second term in (6.3), we again use co-area formula as follows:
$\int_{\Omega^{\prime}_{\varepsilon}}\mathbb{P}_{x}\left(\sup_{0\leqslant s\leqslant t}|X_{s}-x|^{1+\kappa}>(1-\eta)\phi(x)\right)dx$
$=\int_{\Omega^{\prime}_{\varepsilon}}\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}|X_{s}|^{1+\kappa}>(1-\eta)\phi(x)\right)dx$
$=\int_{0}^{\varepsilon}\int_{\partial\Omega_{v}}\frac{\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}|X_{s}|^{1+\kappa}>(1-\eta)v\right)}{|\nabla\phi(x)|}\mathcal{H}^{d-1}(dx)dv$
$\leqslant\sup_{0\leqslant v\leqslant\varepsilon}\mathcal{H}^{d-1}(\partial\Omega_{v})\frac{1}{(1-\eta)m_{\varepsilon}}\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}|X_{s}|^{1+\kappa}\wedge(1-\eta)\varepsilon\right]$
(6.6) $\leqslant 2\sup_{0\leqslant v\leqslant\varepsilon}\mathcal{H}^{d-1}(\partial\Omega_{v})\frac{(1-\eta)^{-\frac{\kappa}{\kappa+1}}}{m_{\varepsilon}}\mu(t)\varepsilon^{\frac{\kappa}{\kappa+1}},$
where in the last inequality we used
$\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}|X_{s}|^{1+\kappa}\wedge(1-\eta)\varepsilon\right]=\mathbb{E}_{0}\left[\left(\sup_{0\leqslant s\leqslant t}|X_{s}|\wedge(1-\eta)^{\frac{1}{1+\kappa}}\varepsilon^{\frac{1}{1+\kappa}}\right)^{\kappa+1}\right]$
$\leqslant((1-\eta)\varepsilon)^{\frac{1}{\kappa+1}}\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}|X_{s}|\wedge(1-\eta)^{\frac{1}{1+\kappa}}\varepsilon^{\frac{1}{1+\kappa}}\right]\leqslant 2\mu(t)((1-\eta)\varepsilon)^{\frac{\kappa}{\kappa+1}}$
for any $\varepsilon\in(0,1)$. Since $m_{\varepsilon}$ is bounded away from 0 as $\varepsilon\to 0$, (6.6) implies
(6.7) $\limsup_{t\to 0}\frac{1}{\mu(t)}\int_{\Omega^{\prime}_{\varepsilon}}\mathbb{P}_{x}\left(\sup_{0\leqslant s\leqslant t}|X_{s}-x|^{1+\kappa}>(1-\eta)\phi(x)\right)dx$
$\leqslant 2\sup_{0\leqslant v\leqslant\varepsilon}\mathcal{H}^{d-1}(\partial\Omega_{v})\frac{(1-\eta)^{-\frac{\kappa}{\kappa+1}}}{m_{\varepsilon}}\varepsilon^{\frac{\kappa}{\kappa+1}}.$
Hence, the proof of the lemma follows after combining (6.5) and (6.7). ∎
Proof of Theorem 6.1.
By Proposition 5.1, it suffices to prove that
$\limsup_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\leqslant\frac{\sup_{\partial\Omega}|\nabla\phi|}{\inf_{\partial\Omega}|\nabla\phi|}\mathrm{Per}(\Omega).$
For any $\varepsilon>0$ we have
(6.8) $\ \ \ \ \limsup_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}=\int_{\Omega}\mathbb{P}_{x}(\tau^{X}_{\Omega}\leqslant t)dx$
$\leqslant\limsup_{t\to 0}\frac{1}{\mu(t)}\int_{\Omega_{\varepsilon}}\mathbb{P}_{x}\left(\inf_{0\leqslant s\leqslant t}\phi(X_{s})\leqslant 0\right)dx$
$+\limsup_{t\to 0}\frac{1}{\mu(t)}\int_{\Omega^{\prime}_{\varepsilon}}\mathbb{P}_{x}\left(\inf_{0\leqslant s\leqslant t}\phi(X_{s})\leqslant 0\right)dx.$
By Lemma 6.3, the first term on the right hand side of (6.8) is zero. Recalling Lemma 6.4, we get
$\ \ \ \ \limsup_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}$
$\leqslant 2\sup_{0\leqslant v\leqslant\varepsilon}\mathcal{H}^{d-1}(\partial\Omega_{v})\frac{M_{\varepsilon}}{\eta m_{\varepsilon}}+2\sup_{0\leqslant v\leqslant\varepsilon}\mathcal{H}^{d-1}(\partial\Omega_{v})\frac{(1-\eta)^{-\frac{\kappa}{\kappa+1}}}{m_{\varepsilon}}\varepsilon^{\frac{\kappa}{\kappa+1}}$
As $|\nabla\phi|$ is bounded in $\Omega^{\prime}_{\varepsilon}$ for sufficiently small $\varepsilon>0$, using Lemma 6.2, we have
$\lim_{\varepsilon\to 0}\sup_{0\leqslant v\leqslant\varepsilon}\mathcal{H}^{d-1}(\partial\Omega_{v})\frac{M_{\varepsilon}}{\eta m_{\varepsilon}}=\frac{1}{\eta}\frac{\sup_{\partial\Omega}|\nabla\phi|}{\inf_{\partial\Omega}|\nabla\phi|}\mathcal{H}^{d-1}(\partial\Omega),$
$\lim_{\varepsilon\to 0}\sup_{0\leqslant v\leqslant\varepsilon}\mathcal{H}^{d-1}(\partial\Omega_{v})\frac{(1-\eta)^{-\frac{\kappa}{\kappa+1}}}{m_{\varepsilon}}\varepsilon^{\frac{\kappa}{\kappa+1}}=0.$
Finally, letting $\eta\to 1$, we conclude that
$\limsup_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\leqslant\frac{\sup_{\partial\Omega}|\nabla\phi|}{\inf_{\partial\Omega}|\nabla\phi|}\mathcal{H}^{d-1}(\partial\Omega)=\frac{\sup_{\partial\Omega}|\nabla\phi|}{\inf_{\partial\Omega}|\nabla\phi|}\mathrm{Per}(\Omega).$
This completes the proof of the theorem. ∎
Concluding proof of Theorem 1.1
The final step of the proof of Theorem 1.1 is to construct a function $\phi$ satisfying conditions of Theorem 6.1 for domains with $C^{1,1}$-boundary.
Lemma 6.5.
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded open set with $C^{1,1}$-boundary. Then there exists a bounded $C^{1,1}$ function $\phi:\mathbb{R}^{d}\to\mathbb{R}$ such that
  1. (1)
    $\Omega=\{x:\phi(x)>0\}$,
  2. (2)
    $|\nabla\phi(x)|=1$ in a neighborhood of $\partial\Omega$.
Proof.
Consider the signed distance function
$\delta(x)=\begin{cases}d(x,\partial\Omega)&\mbox{if $x\in\Omega$}\\ -d(x,\partial\Omega)&\mbox{if $x\in\Omega^{c}$}.\end{cases}$
Since $\Omega$ has $C^{1,1}$-boundary, there exists $r>0$ such that $\delta$ is $C^{1,1}$ in $\Omega^{\prime}_{r}:=\{x\in\mathbb{R}^{d}:|\delta(x)|<r\}$ and $|\nabla\delta|=1$ in $\Omega^{\prime}_{r}$. Define a $C^{1,1}$ function $\varphi:\mathbb{R}\to\mathbb{R}$ such that $\varphi(x)=x$ when $|x|<r/2$ and equals $r$ when $|x|>r$ and consider $\phi=\varphi\circ\delta$. Clearly, $\phi$ is $C^{1,1}$ and $\phi(x)>0$ if and only if $x\in\Omega$. Also, for any $x\in\Omega^{\prime}_{r/2}$, $|\nabla\phi(x)|=|\varphi^{\prime}(\delta(x))\nabla\delta(x)|=1$. This completes the proof of the lemma. ∎
Remark 6.6.
The $C^{1,1}$-regularity assumption on $\partial\Omega$ cannot be weakened. From [8, Corollary 6] it is known that the differentiability of $\delta$ in a neighborhood of $\partial\Omega$ implies that $\partial\Omega$ is $C^{1,1}$.
We note that $\phi$ in Lemma 6.5 satisfies the conditions of Theorem 6.1 with $\kappa=1$, and $\Omega=\phi^{-1}(0,\infty)$. Also, $|\nabla\phi|=1$ on $\partial\Omega$. Therefore, Theorem 1.1 follows from Theorem 6.1. ∎
Proof of Corollary 1.3.
For any $\varepsilon>0$, using Markov inequality we obtain
$\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>\varepsilon\right)\leqslant\frac{\mathbb{E}_{0}\left[\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\right)^{\alpha}\right]}{\varepsilon^{\alpha}}.$
Also, from the given conditions and noting that $\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\geqslant 0$ if $X_{0}=0$, we have
$0\leqslant 1-\frac{\mu(t)}{m(t)}\leqslant\frac{\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\mathbbm{1}\left\{\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>1\right\}\right]}{m(t)}\leqslant\frac{\mathbb{E}_{0}\left[\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\right)^{\alpha}\right]}{m(t)},$
and the right hand side goes to 0 as $t\to 0$. Hence, $\mu(t)/m(t)\to 1$ as $t\to 0$ and Assumption 1.1 is satisfied. The proof is concluded by applying Theorem 1.1. ∎

### v3

@@
 Abstract.
-We provide a general approach for proving small time asymptotic of spectral heat content for any translation invariant isotropic process satisfying negligible tail probability condition.
-As a consequence, we recover several existing results in the context of Lévy processes and Gaussian processes, and provide spectral heat content asymptotic for a class of time-changed Brownian motions.
+The spectral heat content of a domain $\Omega\subset\mathbb{R}^{d}$ corresponding to a $d$-dimensional stochastic process $X=(X_{t})_{t\geqslant 0}$ is defined as
+$Q^{X}_{\Omega}(t)=\int_{\mathbb{R}^{d}}\mathbb{P}_{x}(\tau^{X}_{\Omega}>t)dx,$
+where $\tau^{X}_{\Omega}$ is the first exit time of $X$ from $\Omega$.
+We provide a novel technique for proving small time asymptotic of spectral heat content for any translation invariant isotropic process satisfying negligible tail probability condition.
+As a consequence, we recover several existing results in the context of Lévy processes and Gaussian processes, and provide spectral heat content asymptotics for a class of $\alpha$-stable Lévy processes time-changed by right inverse of positive, increasing, self-similar Markov processes.
+The latter has connection to some Cauchy problems that are non-local in both time and space.
 Key words and phrases:
-spectral heat content; isotropic processes; asymptotic behavior
+spectral heat content; isotropic processes; self-similarity; asymptotic behavior
 2020 Mathematics Subject Classification:
-60J45; 60G51; 60G15
+Primary: 60J45; 60G07; 60G15; 60G51. Secondary: 45K05
 1. Introduction and main result
-Let $X=(X_{t})_{t\geqslant 0}$ be a stochastic process on $\mathbb{R}^{d}$ with càdlàg paths and $\mathbb{P}_{x}$ denotes the probability measure when $X_{0}=x$.
-For any bounded open set $\Omega\subset\mathbb{R}^{d}$, the spectral heat content of $X$ in the domain $\Omega$ is defined as
+1.1. Main result
+Let $X=(X_{t})_{t\geqslant 0}$ be a stochastic process on $\mathbb{R}^{d}$ with càdlàg paths, and let $\mathbb{P}_{x}$ denotes the probability measure when $X_{0}=x$.
+For any bounded open set $\Omega\subset\mathbb{R}^{d}$, the spectral heat content (SHC) of $X$ in the domain $\Omega$ is defined as
 $Q^{X}_{\Omega}(t)=\int_{\Omega}\mathbb{P}_{x}(\tau^{X}_{\Omega}>t)dx,$
 where $\tau^{X}_{\Omega}=\inf\{t\geqslant 0:X_{t}\notin\Omega\}$. Since $X$ has càdlàg paths, it follows that
-$\lim_{t\to 0}Q^{X}_{\Omega}(t)=\mathrm{Vol}(\Omega),$
+$\lim_{t\downarrow 0}Q^{X}_{\Omega}(t)=\mathrm{Vol}(\Omega),$
 where $\mathrm{Vol}(\Omega)$ denotes the volume of $\Omega$.
@@
 From right continuity of the sample paths we have
-$\lim_{t\to 0}\mu(t)=0.$
-Throughout the article, we assume the negligible tail probability condition.
+$\lim_{t\downarrow 0}\mu(t)=0.$
+Throughout the article, we assume the negligibility of the tail probability in the following sense.
 Assumption 1.1.
 For any $\varepsilon>0$,
-$\lim_{t\to 0}\frac{1}{\mu(t)}\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>\varepsilon\right)=0.$
-The goal of this article is to prove the following asymptotic of the spectral heat content of $X$.
+$\lim_{t\downarrow 0}\frac{1}{\mu(t)}\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>\varepsilon\right)=0.$
+The aim of this paper is to prove the following asymptotic behavior of the spectral heat content of $X$.
 Theorem 1.1.
-Let $\Omega$ be bounded open subset of $\mathbb{R}^{d}$ with $C^{1,1}$-boundary, and $X=(X_{t})_{t\geqslant 0}$ be a translation invariant isotropic stochastic process satisfying Assumption 1.1. Then,
-(1.2) $\lim_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}=\mathrm{Per}(\Omega),$
+Let $\Omega$ be a bounded open subset of $\mathbb{R}^{d}$ with $C^{1,1}$ boundary, and $X=(X_{t})_{t\geqslant 0}$ be a translation invariant isotropic stochastic process satisfying Assumption 1.1.
+Then,
+(1.2) $\lim_{t\downarrow 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}=\mathrm{Per}(\Omega),$
 where $\mathrm{Per}(\Omega)$ denotes the perimeter of $\Omega$.
 Remark 1.2.
-The $C^{1,1}$ regularity condition on the boundary of $\Omega$ inherits from the differentiability of the signed distance function from the boundary, see Remark 6.6 for details.
-Corollary 1.3.
-Let $\Omega$ be a bounded, open set with $C^{1,1}$-boundary and $X=(X_{t})_{t\geqslant 0}$ be a translation invariant isotropic process such that
-$\mathbb{E}_{0}\left[\left(\sup_{0\leqslant s\leqslant T}X^{(1)}_{s}\right)^{\alpha}\right]<\infty$
-for some $T>0$ and $\alpha>1$, and
-$\lim_{t\to 0}\frac{\mathbb{E}_{0}\left[\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\right)^{\alpha}\right]}{\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\right]}=0.$
+The $C^{1,1}$ regularity condition on the boundary of $\Omega$ inherits from the differentiability of the signed distance function from the boundary, which is crucial in our proof of the above theorem.
+We refer to Remark 6.6 for details.
+Assumption 1.1 is satisfied by a wide range of isotropic processes.
+The next result provides sufficient conditions for Assumption 1.1 to hold in terms of the moments of the supremum process $\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}$.
+For the ease of writing, we introduce the following notation to be used throughout the rest.
+Notation 1.1.
+For any stochastic process $Z=(Z_{t})_{t\geqslant 0}$ on $\mathbb{R}$, we denote the running supremum of $Z$ by $\overline{Z}$, that is,
+$\overline{Z}_{t}=\sup\{Z_{s}:0\leqslant s\leqslant t\}.$
+Proposition 1.3.
+Let $\Omega$ be a bounded, open set with $C^{1,1}$ boundary and $X=(X_{t})_{t\geqslant 0}$ be a translation invariant isotropic process such that
+$\mathbb{E}_{0}\left[\left(\overline{X}^{(1)}_{T}\right)^{p}\right]<\infty$
+for some $T>0$ and $p>1$, and
+$\lim_{t\downarrow 0}\frac{\mathbb{E}_{0}\left[\left(\overline{X}^{(1)}_{t}\right)^{p}\right]}{\mathbb{E}_{0}\left[\overline{X}^{(1)}_{t}\right]}=0.$
 Then,
-$\lim_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{m(t)}=\mathrm{Per}(\Omega),$
-where $m(t)=\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\right$].
-When $X=B$, the standard Brownian motion on $\mathbb{R}^{d}$, the spectral heat content can be realized in terms of the Dirichlet heat equation. More precisely, if $u(t,x)$ solves the PDE
-$\frac{\partial u}{\partial t}u(x,t)$ $=\frac{1}{2}\Delta u(x,t)\quad\mbox{in $\Omega\times(0,\infty)$}$
+$\lim_{t\downarrow 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mathbb{E}_{0}\left[\overline{X}^{(1)}_{t}\right]}=\mathrm{Per}(\Omega).$
+1.2. Previous wokrs
+The spectral heat content in the classical sense is defined via heat equation with Dirichlet boundary condition on $\Omega$. More precisely, if $u(t,x)$ solves the PDE
+$\frac{\partial}{\partial t}u(x,t)$ $=\Delta_{x}u(x,t)\quad\mbox{in $\Omega\times(0,\infty)$}$
 $u(x,t)$ $=0\quad\mbox{for $(x,t)\in\partial\Omega\times(0,\infty)$}$
 $u(x,0)$ $=1\quad\mbox{for $x\in\Omega$},$
-then the spectral heat content on $\Omega$ for $B$ is given by
-$Q^{B}_{\Omega}(t)=\int_{\Omega}u(x,t)dx.$
-Small time asymptotic of $Q^{B}_{\Omega}(t)$ was first proved by Berg, see [12, Theorem 6.2], for domains with $R$-smooth boundary, and subsequently, a second order expansion of $Q^{B}_{\Omega}(t)$ near zero was proved in [14, Theorem 1.2] for domains with $C^{3}$-boundary.
-We also refer to [13] for small time asymptotic of $Q^{B}_{\Omega}$ when $B$ is the Brownian motion on a Riemannian manifold and $\Omega$ has smooth boundary.
-The spectral heat content asymptotic for symmetric $\alpha$-stable was obtained in [9], and it was later generalized to isotropic Lévy processes in [4], and to symmetric Lévy processes in a recent article [6].
-We also refer to [5] for the spectral heat content asymptotic of Gaussian processes in $\mathbb{R}^{d}$.
-In the present article, we provide a general approach for spectral heat content asymptotic that works for any translation invariant isotropic process satisfying Assumption 1.1.
+then the spectral heat content of $\Omega$ is given by
+$Q_{\Omega}(t)=\int_{\Omega}u(x,t)dx.$
+From standard theory of Markov processes, one can write $Q_{\Omega}(t)=Q^{B}_{\Omega}(t)$, where $B$ is the standard Brownian motion on $\mathbb{R}^{d}$.
+Small time asymptotics of $Q^{B}_{\Omega}(t)$ was first proved by van den Berg and Davies, see [25, Theorem 6.2], for domains with $R$-smooth boundary, and subsequently, van den Berg and LeGall [27, Theorem 1.2] obtained a second order expansion of $Q^{B}_{\Omega}(t)$ for domains with $C^{3}$ boundary as follows:
+$Q^{B}_{\Omega}(t)=\mathrm{Vol}(\Omega)-\sqrt{\frac{2t}{\pi}}\sigma(\partial\Omega)+\frac{t}{4}\int_{\partial\Omega}H_{\Omega}d\sigma+\mathrm{O}(t^{3/2}),$
+where $\sigma$ is the surface measure on $\partial\Omega$, and $H_{\Omega}$ is the mean curvature at the boundary.
+The above formula was further generalized to Riemannian manifolds by van den Berg and Gilkey [26] for domains with smooth boundary.
+Using PDE techniques, Savo [23] proved an asymptotic expansion of $Q_{\Omega}(t)$ of arbitrary order for any domain with smooth boundary in a Riemannian manifold.
+While the asymptotics for Brownian motion is known for several years, results for jump processes have been obtained quite recently.
+In the one-dimensional case, SHC asymptotic for symmetric $\alpha$-stable process was obtained by Valverde [2], and later generalized to isotropic $\alpha$-stable processes on $\mathbb{R}^{d}$ by Park and Song [18].
+We refer to [10] for SHC of isotropic Lévy processes with regularly varying Lévy-Khintchine exponents, and to a recent article [13] for symmetric Lévy processes (may not be isotropic) with some conditions on the Lévy-Khintchine exponent.
+When the Lévy process has bounded variation, the small time asymptotic of SHC was proved in [7], and in this case, one recovers the fractional perimeter in the limit.
+Beyond Lévy processes, we refer to [11, 9, 8] for separate treatments of SHC asymptotics of time-changed killed Brownian motion, time-changed killed $\alpha$-stable processes, and Gaussian processes in $\mathbb{R}^{d}$.
+1.3. Our contribution
+In the present article, we propose a novel approach based on simple facts from geometric measure theory for proving SHC asymptotic for any translation invariant isotropic process satisfying Assumption 1.1.
 This assumption holds true for any isotropic Lévy process with regularly varying Lévy-Khintchine exponent, and Gaussian processes.
-Moreover, in Section 2.3, we provide examples of isotropic processes which are neither Markov nor Gaussian and Theorem 1.1 holds for such processes.
-Our method also extends to bounded domains with boundary being level sets of a $C^{1,\kappa}$ functions for $0<\kappa<1$.
-In all the aforementioned references [12, 13, 14, 9, 4, 5], the assumption of $C^{1,1}$-regularity is crucial as the proofs rely heavily on the uniform interior and exterior ball condition of the domain.
-In fact, from [7, Corollary 2], the uniform interior and exterior ball condition is equivalent to the $C^{1,1}$ smoothness of the boundary.
-For $C^{1,\kappa}$-boundary with $0<\kappa<1$, the uniform ball condition does not hold.
-Theorem 6.1 provides upper and lower bound of the spectral heat content for domains with boundaries given by level sets of $C^{1,\kappa}$ functions with $0<\kappa<1$.
+Moreover, in Section 2.3, we provide a wide range of examples of time-changed isotropic $\alpha$-stable processes that satisfy Assumption 1.1.
+As a consequence, we recover the results in [18, 8, 9, 10], and in Section 2.3.2, we provide examples where the SHC is related to some integro-partial differential equations that are non-local in both time and space.
+Another important aspect of our method is that it requires very minimal assumption on the isotropic process in terms of the tail probability behavior of the supremum of the one dimensional marginals.
+This observation led us to adapt this technique for heat content of fractional sub-Laplacian on Carnot groups and we refer the interested reader to [21].
+Our techniques can be applied to bounded domains with boundary being level sets of a $C^{1,\kappa}$ functions for $0<\kappa<1$.
+In all the aforementioned references [25, 26, 27, 18, 10, 11], the assumption of $C^{1,1}$ regularity is crucial as the proofs rely heavily on the uniform interior and exterior ball condition of the domain.
+In fact, from [14, Corollary 2], the uniform interior and exterior ball condition is equivalent to the $C^{1,1}$ smoothness of the boundary.
+For $C^{1,\kappa}$ boundary with $0<\kappa<1$, the uniform ball condition does not hold.
+In Theorem 6.1 we provide upper and lower bound of the SHC for domains with boundaries given by level sets of $C^{1,\kappa}$ functions with $0<\kappa<1$.
 The question of validity of (1.2) for such domains still remains open.
-We propose a novel approach by considering a functional version of spectral heat content. For any nonnegative integrable function $f:\mathbb{R}^{d}\to\mathbb{R}$ we define
-(1.3) $Q^{X}_{f}(t)=\int_{\mathbb{R}^{d}}\mathbb{E}_{x}\left[\inf_{0\leqslant s\leqslant t}f(X_{s})\right]dx,$
+Let us now briefly describe the main idea behind our approach to prove Theorem 1.1. For any nonnegative integrable function $f:\mathbb{R}^{d}\to\mathbb{R}$ we define
+(1.3) $Q^{X}_{f}(t)=\int_{\mathbb{R}^{d}}\mathbb{E}_{x}\left[\inf_{0\leqslant s\leqslant t}f(X_{s})\right]dx.$
 In particular, when $f=\mathbbm{1}_{\Omega}$, $Q^{X}_{f}(t)=Q^{X}_{\Omega}(t)$ for all $t\geqslant 0$.
-First, we obtain asymptotic formula for $Q^{X}_{f}$ when $f$ is a compactly supported smooth function, and the lower bound of the limit in (1.2) is proved using an approximation argument.
-It turns out that the lower bound of (1.2) holds for any bounded, open subset with finite perimeter.
-The main idea behind proving the upper bound of the limit in (1.2) stems from the observation that any bounded domain with $C^{1,1}$-boundary can be expressed as $\phi^{-1}(0,\infty)$ for some bounded $C^{1,1}$ function $\phi$ with $|\nabla\phi|$ being constant on the boundary.
-The rest of the paper is organized as follows: in Section 2 we provide examples including isotropic Lévy processes, Gaussian processes, and a class of time-changed Brownian motions for which spectral heat content asymptotic follow directly from our main result.
+Using Taylor expansion of $f$, we first obtain the asymptotic of $Q^{X}_{f}$ when $f$ is a compactly supported smooth function, and the lower bound of the limit in (1.2) is proved using an approximation argument.
+It turns out that the lower bound of (1.2) holds for any bounded, open subset with finite perimeter, see Proposition 5.1.
+The key idea to derive the upper bound of the limit in (1.2) stems from the observation that any bounded domain with $C^{1,1}$ boundary can be expressed as $\phi^{-1}(0,\infty)$ for some bounded $C^{1,1}$ function $\phi$ with $|\nabla\phi|$ being constant on the boundary.
+The rest of the paper is organized as follows: in Section 2 we provide examples including isotropic Lévy processes, Gaussian processes, and a class of time-changed $\alpha$-stable isotropic Lévy processes for which spectral heat content asymptotics follow directly from our main result.
 Section 3 provides some preliminaries about functions of bounded variation and regularity of boundary.
@@
 Let $X=(X_{t})_{t\geqslant 0}$ be an isotropic Lévy process with Lévy-Khintchine exponent $\Psi(\xi)=\psi(|\xi|)$ for any $\xi\in\mathbb{R}^{d}$, and assume that $\psi:[0,\infty)\to[0,\infty)$ is regularly varying with exponent $\beta\in[1,2]$.
-Then, by [10, Equation (3.2)], for any $\varepsilon>0$,
-$\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>\varepsilon\right)\leqslant c_{\varepsilon}t$
+Then, by [20, Equation (3.2)], for any $\varepsilon>0$,
+$\mathbb{P}_{0}\left(\overline{X}^{(1)}_{t}>\varepsilon\right)\leqslant c_{\varepsilon}t$
 for some constant $c_{\varepsilon}>0$.
-On the other hand, using the bounds in [6, Equation (3.8)], we have $\lim_{t\to 0}\frac{t}{\mu(t)}=0$, which shows that Assumption 1.1 holds.
-Therefore, Theorem 1.1 holds for $X$, and we recover [6, Theorem 2.1(1)] for isotropic Lévy processes.
-In particular, when $X_{t}$ is the symmetric $\alpha$-stable process, as $t\to 0$ we have
-$\mu(t)\sim\begin{cases}t^{1/\alpha}\mathbb{E}[\sup_{0\leqslant s\leqslant 1}X^{(1)}_{s}]&\mbox{if $1<\alpha\leqslant 2$}\\ \frac{1}{\pi}t\log(1/t)&\mbox{if $\alpha=1$}.\end{cases}$
+On the other hand, using the bounds in [13, Equation (3.8)], we have $\lim_{t\downarrow 0}\frac{t}{\mu(t)}=0$, which shows that Assumption 1.1 holds.
+Therefore, Theorem 1.1 holds for $X$, and we recover [13, Theorem 2.1(1)] for isotropic Lévy processes.
+In particular, when $X$ is the isotropic $\alpha$-stable Lévy process, that is, $\Psi(\xi)=|\xi|^{\alpha}$, as $t\downarrow 0$ we have
+$\mu(t)\sim\begin{cases}t^{1/\alpha}\mathbb{E}[\overline{X}^{(1)}_{t}]&\mbox{if $1<\alpha\leqslant 2$},\\ \frac{1}{\pi}t\log(1/t)&\mbox{if $\alpha=1$}.\end{cases}$
 For $\alpha\in(1,2]$, the asymptotic behavior of $\mu(t)$ above follows directly from the self-similarity of $\alpha$-stable process, and for $\alpha=1$, we refer to [2, Proposition 4.3(i)].
-This recovers [9, Theorem 1.1] when $\alpha\geqslant 1$.
+This recovers [18, Theorem 1.1] when $\alpha\geqslant 1$.
 2.2. Gaussian processes
@@
 For any $t>0$, let
-$m(t)=\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\right],\quad\sigma^{2}(t)=\sup_{0\leqslant s\leqslant t}\mathrm{Var}\left[(X^{(1)}_{s})^{2}\right].$
-It is known that $m(t)<\infty$ for all $t\geqslant 0$. By Borell-TIS inequality, see [3, Theorem 2.1.1], for any $x>0$
-(2.1) $\mathbb{P}_{0}\left(\left|\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}-m(t)\right|>x\right)$ $\leqslant 2e^{-x^{2}/2\sigma^{2}(t)},$
+$m(t)=\mathbb{E}_{0}\left[\overline{X}^{(1)}_{t}\right],\quad\sigma(t)^{2}=\sup_{0\leqslant s\leqslant t}\mathrm{Var}\left[(X^{(1)}_{s})^{2}\right].$
+We assume that $m(t)<\infty$ for all $t\in[0,1]$, which is known to be equivalent to the condition that $\mathbb{P}_{0}(\overline{X}^{(1)}_{1}<\infty)=1$.
+By Borell-TIS inequality, see [3, Theorem 2.1.1], for any $x>0$
+(2.1) $\mathbb{P}_{0}\left(\left|\overline{X}^{(1)}_{t}-m(t)\right|>x\right)$ $\leqslant 2e^{-x^{2}/2\sigma(t)^{2}},$
 and hence
-$\mathrm{Var}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\right]\leqslant\int_{0}^{\infty}4xe^{-x^{2}/2\sigma^{2}(t)}=4\sigma^{2}(t).$
-Therefore, $\mathbb{E}_{0}\left[\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\right)^{2}\right]\leqslant m(t)^{2}+4\sigma^{2}(t)$.
-From [4, Equation (29)] it is known that $\lim_{t\to 0}{\sigma^{2}(t)}/{m(t)}=0$.
-Applying Corollary 1.3, we obtain the following result.
+$\mathrm{Var}\left[\overline{X}^{(1)}_{t}\right]\leqslant\int_{0}^{\infty}4xe^{-x^{2}/2\sigma(t)^{2}}=4\sigma(t)^{2}.$
+Therefore, $\mathbb{E}_{0}\left[\left(\overline{X}^{(1)}_{t}\right)^{2}\right]\leqslant m(t)^{2}+4\sigma(t)^{2}$.
+From [10, Equation (29)] it is known that $\lim_{t\downarrow 0}{\sigma(t)^{2}}/{m(t)}=0$.
+Applying Proposition 1.3, we obtain the following result.
 Proposition 2.1.
-If $\Omega$ is a bounded open set with $C^{1,1}$-boundary then,
-(2.2) $\lim_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{m(t)}=\mathrm{Per}(\Omega).$
+If $\Omega$ is a bounded open set with $C^{1,1}$ boundary then,
+(2.2) $\lim_{t\downarrow 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{m(t)}=\mathrm{Per}(\Omega).$
 Remark $2.$
-In [4, Theorem 4.4] under the condition that the rescaled process $Y^{(t)}_{u}=\mu^{-1}_{t}X^{(1)}_{tu}$ converges weakly to a Gaussian process $(Y_{u})_{u\geqslant 0}$ as $t\downarrow 0$ in Skorokhod $J_{1}$ topology, it is proved that
-$\lim_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{m(t)}=\mathbb{E}\left[\sup_{0\leqslant u\leqslant 1}Y_{u}\right]\mathrm{Per}(\Omega),$
-with $\mathbb{E}[\sup_{0\leqslant u\leqslant 1}Y_{u}]\leqslant 1$. Proposition 2.1 strengthens [4, Theorem 4.4].
-2.3. Time changed Brownian motion
-Consider the $d$-dimensional standard Brownian motion $B=(B_{t})_{t\geqslant 0}$ and a nonnegative, increasing continuous process $U=(U_{t})_{t\geqslant 0}$ which is independent of $B$ and $U_{0}=0$.
-Define $X_{t}=B_{U_{t}}$.
+In [10, Theorem 4.4] under the condition that the rescaled process $Y^{(t)}_{u}=\mu^{-1}_{t}X^{(1)}_{tu}$ converges weakly to a Gaussian process $(Y_{u})_{u\geqslant 0}$ as $t\downarrow 0$ in Skorokhod $J_{1}$ topology, it is proved that
+$\lim_{t\downarrow 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{m(t)}=\mathbb{E}\left[\overline{Y}_{\!1}\right]\mathrm{Per}(\Omega),$
+with $\mathbb{E}[\overline{Y}_{\!1}]\leqslant 1$. Proposition 2.1 does not require the convergence of the rescaled Gaussian process, and hence it strengthens [10, Theorem 4.4].
+2.3. Time-changed $\alpha$-stable process
+Consider the $d$-dimensional $\alpha$-stable isotropic Lévy process $Y=(Y_{t})_{t\geqslant 0}$ with $\alpha\in(1,2]$, and a nonnegative, increasing process $U=(U_{t})_{t\geqslant 0}$ with càdlàg paths which is independent of $Y$ and $U_{0}=0$.
+Define $X_{t}=Y_{U_{t}}$.
 Then, $X=(X_{t})_{t\geqslant 0}$ is translation invariant and isotropic.
-Let us assume that for some $\alpha>1/2$,
-(2.3) $\mathbb{E}[U^{\alpha}_{t}]<\infty\quad\forall t\geqslant 0,\quad\lim_{t\to 0}\frac{\mathbb{E}[U^{\alpha}_{t}]}{\mathbb{E}[\sqrt{U_{t}}]}=0.$
+Let us assume that for some $p>1$,
+(2.3) $\mathbb{E}[U^{p/\alpha}_{t}]<\infty\quad\text{for all $t\geqslant 0$},\quad\lim_{t\downarrow 0}\frac{\mathbb{E}[U^{p/\alpha}_{t}]}{\mathbb{E}[U^{1/\alpha}_{t}]}=0.$
 Then we have the following result.
 Proposition 2.3.
-For any bounded open set $\Omega$ with $C^{1,1}$-boundary we have
-$\lim_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mathbb{E}[\sqrt{U_{t}}]}=\sqrt{\frac{2}{\pi}}\mathrm{Per}(\Omega).$
+For any bounded open set $\Omega$ with $C^{1,1}$ boundary we have
+(2.4) $\lim_{t\downarrow 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{m(t)}=\mathrm{Per}(\Omega),$
+where $m(t)=\mathbb{E}_{0}\left[\overline{X}^{(1)}_{t}\right$]. In particular when $U_{t}$ has continuous paths, we have $m(t)=\mathbb{E}[\overline{Y}^{(1)}_{1}]\mathbb{E}[U^{1/\alpha}_{t}]$.
+Remark 2.4.
+In the context of time-changed Lévy processes, SHC of the time-changed process and the killed time-changed process are different in general.
+This happens as the paths of $U$ are discontinuous in general.
+However when $U$ has continuous paths, above two notions of SHC coincide since
+$\mathbb{P}_{x}\left(\tau^{X}_{\Omega}>t\right)=\mathbb{P}_{x}\left(\tau^{Y}_{\Omega}>U_{t}\right)$
+for all $x\in\Omega$ and $t\geqslant 0$. We refer to [9, p. 5] (see also [24, p. 580]) for detailed discussions on the comparison between killed time-changed and time-changed killed processes.
 Proof.
-Since $U_{t}$ is increasing and continuous, for any $t\geqslant 0$ we have
-$\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}B^{(1)}_{U_{t}}\right]=\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant U_{t}}B^{(1)}_{s}\right]=\int_{0}^{\infty}\mathbb{E}_{0}\left[\sup_{0\leqslant u\leqslant s}B^{(1)}_{u}\right]\mathbb{P}(U_{t}\in ds).$
-Recalling that for any $p>0$,
-$\mathbb{E}_{0}\left[\left(\sup_{0\leqslant s\leqslant t}B^{(1)}_{s}\right)^{p}\right]=\frac{2^{p/2}t^{p/2}\Gamma(\frac{p+1}{2})}{\sqrt{\pi}}$
-we get
-(2.4) $\mathbb{E}_{0}\left[\left(\sup_{0\leqslant s\leqslant t}B^{(1)}_{U_{s}}\right)^{p}\right]=\frac{2^{p/2}\Gamma(\frac{p+1}{2})}{\sqrt{\pi}}\mathbb{E}[U^{p/2}_{t}].$
-Therefore, (2.3) implies that the conditions of Corollary 1.3 hold. This concludes the proof. ∎
-As an example of $U_{t}$, let us consider an $n$-dimensional standard Brownian motion $W=(W_{t})_{t\geqslant 0}$ independent of $B$ and define the integrated squared radial process
-$R_{t}=\int_{0}^{t}|W_{s}|^{2}ds.$
-Then, the one-dimensional marginal distributions of $X_{t}=B_{R_{t}}$ coincide with the Lévy stochastic area, see [11, Equation (3.5)].
-Also, $X$ is neither Markov nor Gaussian.
-For any $t\geqslant 0$,
-$\mathbb{E}[R_{t}]=\int_{0}^{t}nsds=\frac{nt^{2}}{2},$
-and by Cauchy-Schwarz inequality $\mathbb{E}[\sqrt{R_{t}}]\leqslant\sqrt{\mathbb{E}[R_{t}]}<\infty$. On the other hand, using Cauchy-Schwarz inequality, we have
-$\sqrt{\frac{R_{t}}{t}}\geqslant\frac{1}{t}\int_{0}^{t}|W_{s}|ds,$
+Since $Y$ is an $\alpha$-stable Lévy process, for any $0\leqslant q<\alpha$, $\mathbb{E}[\sup_{0\leqslant s\leqslant t}|Y_{t}|^{q}]<\infty$, see for instance [22, Theorem 25.18].
+By $1/\alpha$–self-similarity of $Y$, we have $\mathbb{E}[(\overline{Y}^{(1)}_{t})^{q}]=t^{q/\alpha}\mathbb{E}[(\overline{Y}^{(1)}_{1})^{q}]$.
+Since $U$ is increasing and right continuous, for any $t\geqslant 0$,
+$\overline{X}^{(1)}_{t}\leqslant\sup_{0\leqslant s\leqslant U_{t}}Y^{(1)}_{s}$,
 which shows that
-$\mathbb{E}[\sqrt{R_{t}}]\geqslant\frac{1}{\sqrt{t}}\mathbb{E}\left[\int_{0}^{t}|W_{s}|ds\right]\geqslant ct$
-for some constant $c>0$. Therefore, $\lim_{t\to 0}\mathbb{E}[R_{t}]/\mathbb{E}[\sqrt{R_{t}}]=0$, and hence the conditions in (2.3) are satisfied by $R_{t}$, which implies the following corollary.
-Corollary 2.4.
-Let the processes $B,R$ be defined as above and $X_{t}=B_{R_{t}}$. Then, for any bounded, open set $\Omega$ with $C^{1,1}$-boundary
-$\lim_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mathbb{E}[\sqrt{R_{t}}]}=\sqrt{\frac{2}{\pi}}\mathrm{Per}(\Omega).$
+$\mathbb{E}\left[\left(\overline{X}^{(1)}_{t}\right)^{q}\right]\leqslant\mathbb{E}_{0}\left[\left(\overline{Y}^{(1)}_{1}\right)^{q}\right]\mathbb{E}\left[U^{q/\alpha}_{t}\right]<\infty$
+for $0\leqslant q<p\wedge\alpha$. On the other hand, using symmetry of the distribution of $X^{(1)}$, we note that for any $u>0$,
+$\mathbb{P}_{0}\left(\overline{X}^{(1)}_{t}\geqslant u\right)\geqslant\frac{1}{2}\mathbb{P}_{0}\left(|X^{(1)}_{t}|\geqslant u\right),$
+which implies that for any $1\leqslant q<p\wedge\alpha$,
+$\mathbb{E}_{0}\left[\left(\overline{X}^{(1)}_{t}\right)^{q}\right]\geqslant\frac{1}{2}\mathbb{E}_{0}\left[|X^{(1)}_{t}|^{q}\right]=\frac{1}{2}\mathbb{E}\left[|Y^{(1)}_{1}|^{q}\right]\mathbb{E}[U^{q/\alpha}_{t}],$
+where the last equality follows from self-similarity of $Y$. Therefore, for any $t\geqslant 0$ and $0\leqslant q<p\wedge\alpha$,
+$\frac{1}{2}\mathbb{E}_{0}\left[|Y^{(1)}_{1}|^{q}\right]\frac{\mathbb{E}\left[U^{q/\alpha}_{t}\right]}{\mathbb{E}[U^{1/\alpha}_{t}]}\leqslant\frac{\mathbb{E}_{0}\left[\left(\overline{X}^{(1)}_{t}\right)^{q}\right]}{\mathbb{E}_{0}\left[\overline{X}^{(1)}_{t}\right]}\leqslant\mathbb{E}_{0}\left[\left(\overline{Y}^{(1)}_{1}\right)^{q}\right]\frac{\mathbb{E}\left[U^{q/\alpha}_{t}\right]}{\mathbb{E}[U^{1/\alpha}_{t}]}$
+In light of Proposition 1.3, it suffices to show that $\mathbb{E}[U^{q/\alpha}_{t}]/\mathbb{E}[U^{1/\alpha}_{t}]\to 0$ whenever $1<q<p\wedge\alpha$, where $p$ is given by (2.3).
+By Hölder inequality, for any $1<q<p$ we have
+$\frac{\mathbb{E}\left[U^{q/\alpha}_{t}\right]}{\mathbb{E}[U^{1/\alpha}_{t}]}\leqslant\left(\frac{\mathbb{E}\left[U^{p/\alpha}_{t}\right]}{\mathbb{E}[U^{1/\alpha}_{t}]}\right)^{\frac{q-1}{p-1}},$
+which shows that the conditions of Proposition 1.3 hold due to (2.3). This concludes the proof of (2.4).
+When $U$ has continuous paths, one has
+$\sup_{0\leqslant s\leqslant t}Y^{(1)}_{U_{s}}=\sup_{0\leqslant s\leqslant U_{t}}Y^{(1)}_{s}\quad\mbox{$\mathbb{P}_{0}$--almost surely}.$
+By independence of $Y$ and $U$, and the self-similarity of $Y$, we conclude that $m(t)=\mathbb{E}[\overline{Y}^{(1)}_{1}]\mathbb{E}[U^{1/\alpha}_{t}]$. ∎
+2.3.1. Time-change by inverse subordinator
+Consider a subordinator (increasing Lévy process on $(0,\infty)$, see Chapter III in Bertoin [4]) $S=(S_{t})_{t\geqslant 0}$ with $S_{0}=0$ such that
+$\mathbb{E}[e^{-\lambda S_{t}}]=e^{-t\varphi(\lambda)}\quad\mbox{for all $\lambda,t>0$},$
+where $\varphi$ is a Bernstein function such that
+(2.5) $\varphi(\lambda)=b\lambda+\int_{0}^{\infty}(1-e^{-\lambda r})\mu(dr)$
+with $b\geqslant 0$, and $\int_{0}^{\infty}\min\{1,r\}\mu(dr)<\infty$. The inverse subordinator $E=(E_{t})_{t\geqslant 0}$ is defined as the right inverse of $(S_{t})_{t\geqslant 0}$, that is,
+$E_{t}=\inf\{s\geqslant 0:S_{s}>t\}.$
+Then, $E$ is also an increasing process, and it is continuous in time if and only if $S$ has strictly increasing sample paths almost surely, which occurs if and only if $\varphi(\infty)=\infty$.
+Proposition 2.5.
+Assume that $\varphi$ is regularly varying at infinity with index $\beta\in(0,1)$ and $X_{t}=Y_{E_{t}}$, where $Y$ is a $d$-dimensional isotropic $\alpha$-stable Lévy process with $\alpha\in(1,2]$ and independent of $S$.
+Then,
+$\lim_{t\downarrow 0}\varphi(1/t)^{\frac{1}{\alpha}}(\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t))=\frac{\Gamma(1+1/\alpha)}{\Gamma(1+\beta/\alpha)}\mathbb{E}_{0}\left[\overline{Y}^{(1)}_{1}\right]\mathrm{Per}(\Omega).$
+Remark 2.6.
+The above proposition has been proved in [9, Theorem 4.4] for $\alpha=2$, and in [8] for $\alpha\in(1,2)$. Below we provide an alternative argument using Proposition 2.3.
+Remark 2.7.
+When $\varphi(\lambda)=\lambda^{\beta}$, $\beta\in(0,1)$ and $\alpha=2$, $Q^{X}_{\Omega}(t)=\int_{\Omega}u_{\beta}(x,t)dx$, where $u_{\beta}$ is the unique solution to the time-fractional Cauchy problem with Dirichlet boundary condition on $\Omega$, see [16] for details.
+Proof.
+When $\varphi$ is regularly varying at $\infty$, $\varphi(\infty)=\infty$. Therefore from the discussion above, $E$ has continuous sample paths. By [9, Proposition 4.2], for any $q>0$ one has
+(2.6) $\mathbb{E}\left[E^{q}_{t}\right]\sim\frac{\Gamma(q+1)}{\Gamma(q\beta+1)}\varphi(1/t)^{-q}\quad\mbox{as $t\downarrow 0$}.$
+Hence, for any $p>1$,
+$\frac{\mathbb{E}[E^{p/\alpha}_{t}]}{\mathbb{E}[E^{1/\alpha}_{t}]}\sim\frac{\Gamma(p/\alpha+1)\Gamma(\frac{\beta}{\alpha}+1)}{\Gamma(1+1/\alpha)\Gamma(p\beta/\alpha+1)}\frac{\varphi(1/t)^{1/\alpha}}{\varphi(1/t)^{p/\alpha}}\quad\mbox{as $t\downarrow 0$},$
+which shows that $\lim_{t\downarrow 0}\mathbb{E}[E^{p/\alpha}_{t}]/\mathbb{E}[E^{1/\alpha}_{t}]=0$, since $\varphi$ is regularly varying at $\infty$.
+Therefore using Proposition 2.3 and (2.6) for $q=1/\alpha$, we conclude the proof.
+∎
+2.3.2. Time-change by any positive, increasing self-similar process
+Let $U=(U_{t})_{t\geqslant 0}$ be any positive, $\beta$–self-similar, increasing process with càdlàg paths such that $\mathbb{E}[U^{p/\alpha}_{1}]<\infty$ for some $p>1$, and $Y$ is same as before.
+Then, $X_{t}=Y_{U_{t}}$ is $\beta/\alpha$–self-similar, that is,
+$(cX_{t})_{t\geqslant 0}\overset{d}{=}(X_{tc^{\frac{\alpha}{\beta}}})_{t\geqslant 0}\quad\mbox{for all $c>0$}.$
+By self-similarity, we also have
+$\mathbb{E}[U^{q/\alpha}_{t}]$ $=t^{q\beta/\alpha}\mathbb{E}[U^{q/\alpha}_{1}]\quad\mbox{for $q\in[0,p]$},$
+$\mathbb{E}\left[\overline{X}^{(1)}_{t}\right$] $=t^{\frac{\beta}{\alpha}}\mathbb{E}\left[\overline{X}^{(1)}_{1}\right$].
+Therefore, by Proposition 2.3 we deduce the following result.
+Proposition 2.8.
+For any bounded domain $\Omega$ with $C^{1,1}$ boundary,
+$\lim_{t\downarrow 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{t^{\frac{\beta}{\alpha}}}=\mathbb{E}\left[\overline{X}^{(1)}_{1}\right]\mathrm{Per}(\Omega).$
+We now provide a specific example of a self-similar increasing process constructed from an increasing self-similar Markov process.
+Let $\xi=(\xi_{t})_{t\geqslant 0}$ be a positive, increasing, self-similar Markov process with self-similarity index $1/\beta$, that is, for all $c>0$,
+$\left(c\xi_{tc^{-\beta}},\mathbb{P}_{x}\right)\overset{d}{=}\left(\xi_{t},\mathbb{P}_{cx}\right)$
+By Lamperti’s theorem, see [12], under the probability law $\mathbb{P}_{x}$ for any $x>0$, one can write $\xi$ as a time-changed subordinator as follows:
+$\xi_{t}=x\exp\left(S_{A_{tx^{-\beta}}}\right),$
+where $S$ is a subordinator and $A_{t}=\inf\{s\geqslant 0:\int_{0}^{s}\exp(\beta S_{r})dr>t\}$.
+Moreover, the above representation gives a bijection between the class of all positive, increasing, self-similar Markov processes and subordinators.
+Let $\varphi$ denote the Bernstein function associated with $S$ defined as in (2.5).
+For simplicity, we exclude the case when $S$ is arithmetic, that is, $b=0$ and $\mu$ is supported on a lattice $r\mathbb{Z}$ for some $r>0$.
+When $\varphi^{\prime}(0+)<\infty$, or equivalently, $\mathbb{E}[S_{1}]<\infty$, Bertoin and Caballero [5] proved that the process $(\xi_{t},\mathbb{P}_{x})$ converges weakly as $x\downarrow 0$, and therefore, one gets an increasing, self-similar process $\xi$ with $\xi_{0}=0$.
+Let $\zeta=(\zeta_{t})_{t\geqslant 0}$ denote the right inverse of $\xi$, that is,
+$\zeta_{t}=\inf\{s>0:\xi_{s}>t\}.$
+Then, $\zeta$ is $\beta$-self-similar, positive, and increasing process but may not be Markovian.
+From [15, Lemma 2.1] it is known that $\mathbb{E}[\zeta^{p}_{1}]<\infty$ for all $p>0$.
+For a $d$-dimensional $\alpha$-stable isotropic Lévy process $Y=(Y_{t})_{t\geqslant 0}$ independent of $\xi$, let us define the time-changed process $X_{t}=Y_{\zeta_{t}}$.
+Then, Proposition 2.8 holds for $X$.
+We note that $\zeta$ has discontinuous paths whenever $\varphi(\infty)<\infty$.
+Since $\varphi^{\prime}(0+)<\infty$, using Fubini theorem and a simple change of variable leads to
+$\varphi(\lambda)=b\lambda+\lambda\int_{0}^{1}r^{\lambda-1}m(r)dr,$
+where $m$ is a non-decreasing function on $(0,1)$ and $\int_{0}^{1}(-\log r\wedge 1)m(r)dr<\infty$. Patie and Srapionyan [19, Definition 3.1] introduced the following non-local operator operator:
+$\partial^{\varphi}_{t}f=t^{1-\beta}bf^{\prime}(t)+t^{-\beta}\int_{0}^{t}f^{\prime}(r)m(r/t)dr.$
+In particular, when $\xi$ is the standard $\beta$-stable subordinator with $\beta\in(0,1)$, one has
+$b=0\quad\text{and}\quad m(r)=\frac{(1-r)^{-\beta}}{\Gamma(1-\beta)}\mathbbm{1}_{\{0<r<1\}},$
+see e.g.
+[6, Example 5], and in this case $\partial^{\varphi}_{t}$ coincides with the Caputo fractional derivative of order $\beta$.
+Let us assume that $\zeta$ has continuous paths, or equivalently, $\varphi(\infty)=\infty$.
+Consider the Cauchy problem
+(2.7) $\partial^{\varphi}_{t}u=-(-\Delta_{\Omega})^{\alpha/2}u$
+$u(x,0)=\mathbbm{1}_{\Omega}(x),$
+where $\Delta_{\Omega}$ is the Dirichlet Laplacian on $\Omega$ and $\alpha\in(1,2]$.
+Since $(-\Delta_{\Omega})^{\alpha/2}$ is self-adjoint and compact in $L^{2}(\Omega)$, by [19, Theorem 4.1], $u_{\varphi}(x,t)=\mathbb{P}_{x}\left(\tau^{Y}_{\Omega}>\zeta_{t}\right)$ is a solution to the above integro-partial differential equation.
+Due to path continuity of $\zeta$, we get $u_{\varphi}(x,t)=\mathbb{P}_{x}\left(\tau^{X}_{\Omega}>t\right)$ for all $x\in\Omega$ and $t\geqslant 0$, which shows that the SHC of $X$ is related to the solution to (2.7), that is,
+$Q^{X}_{\Omega}(t)=\int_{\Omega}u_{\varphi}(x,t)dx.$
+In particular, when $\alpha=2$, $u_{\varphi}$ solves the following Cauchy problem with Dirichlet boundary conditions:
+$\partial^{\varphi}_{t}u(x,t)$ $=\Delta_{x}u(x,t)\quad\mbox{in $\Omega\times(0,\infty)$}$
+$u(x,t)$ $=0\quad\mbox{for $(x,t)\in\partial\Omega\times(0,\infty)$}$
+$u(x,0)$ $=1\quad\mbox{for $x\in\Omega$}.$
 3. Preliminaries: Perimeter and regularity of the boundary
@@
 for all $x,y\in\mathbb{R}^{d}$.
-The domain $\Omega\subset\mathbb{R}^{d}$ has $C^{1,\kappa}$-boundary if for every $x\in\partial\Omega$, there exists a neighborhood $V_{x}$ and a $C^{1,\kappa}$ function $\gamma:\mathbb{R}^{d-1}\to\mathbb{R}$ such that
+The domain $\Omega\subset\mathbb{R}^{d}$ has $C^{1,\kappa}$ boundary if for every $x\in\partial\Omega$, there exists a neighborhood $V_{x}$ and a $C^{1,\kappa}$ function $\gamma:\mathbb{R}^{d-1}\to\mathbb{R}$ such that
 $\Omega\cap V_{x}=\{(y_{1},\ldots,y_{d}):y_{d}>\gamma(y_{1},\ldots,y_{d-1})\}\cap V_{x}.$
-If $\Omega$ has $C^{1}$-boundary, from Stoke’s theorem it follows that $\mathrm{Per}(\Omega)=\mathcal{H}^{d-1}(\partial\Omega)$, the (d1)(d-1)-dimensional Hausdorff measure of $\partial\Omega$.
+If $\Omega$ has $C^{1}$ boundary, from Stoke’s theorem it follows that $\mathrm{Per}(\Omega)=\mathcal{H}^{d-1}(\partial\Omega)$, the (d1)(d-1)-dimensional Hausdorff measure of $\partial\Omega$.
 4. The minimum functional
@@
 $Q^{X}_{f}(t)=\int_{\mathbb{R}^{d}}\mathbb{E}_{x}\left[\inf_{0\leqslant s\leqslant t}f(X_{s})\right]dx.$
-In particular, when $f=\mathbbm{1}_{\Omega}$, $Q^{X}_{f}(t)=Q^{X}_{\Omega}(t)$.
+In particular, $Q^{X}_{\mathbbm{1}_{\Omega}}(t)=Q^{X}_{\Omega}(t)$ for any open set $\Omega$.
 Proof.
@@
 $\mathcal{R}^{X}_{f}(t):=\int_{\mathbb{R}^{d}}f(x)dx-Q^{X}_{f}(t).$
-as $t\to 0$. Trivially, when $f=\mathbbm{1}_{\Omega}$, $\mathcal{R}^{X}_{f}(t)=\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)$.
+as $t\downarrow 0$. We note that $\mathcal{R}^{X}_{\mathbbm{1}_{\Omega}}(t)=\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)$.
 Theorem 4.2.
 Let $X=(X_{t})_{t\geqslant 0}$ be a translation invariant, isotropic stochastic process on $\mathbb{R}^{d}$ satisfying Assumption 1.1. Then for any nonnegative $f\in C^{1,\kappa}_{c}(\mathbb{R}^{d})$,
-$\lim_{t\to 0}\frac{\mathcal{R}^{X}_{f}(t)}{\mu(t)}=\int_{\mathbb{R}^{d}}|\nabla f(x)|dx,$
+$\lim_{t\downarrow 0}\frac{\mathcal{R}^{X}_{f}(t)}{\mu(t)}=\int_{\mathbb{R}^{d}}|\nabla f(x)|dx,$
 where $\mu(t)$ is defined in (1.1). In particular, the above limit holds for all compactly supported smooth functions.
@@
 For any $\varepsilon>0$,
-(4.1) $\lim_{t\to 0}\frac{1}{\mu(t)}\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\wedge\varepsilon\right]=1$
-(4.2) $\lim_{t\to 0}\frac{1}{\mu(t)}\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\mathbbm{1}\left\{\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\leqslant\varepsilon\right\}\right]=1$
-(4.3) $\lim_{t\to 0}\frac{1}{\mu(t)}\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\mathbbm{1}\left\{\sup_{0\leqslant s\leqslant t}|X_{s}|\leqslant\varepsilon\right\}\right]=1$
-(4.4) $\limsup_{t\to 0}\frac{1}{\mu(t)}\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}|X_{s}|\mathbbm{1}\left\{\sup_{0\leqslant s\leqslant t}|X_{s}|\leqslant\varepsilon\right\}\right]\leqslant 2d$
+(4.1) $\lim_{t\downarrow 0}\frac{1}{\mu(t)}\mathbb{E}_{0}\left[\overline{X}^{(1)}_{t}\wedge\varepsilon\right]=1$
+(4.2) $\lim_{t\downarrow 0}\frac{1}{\mu(t)}\mathbb{E}_{0}\left[\overline{X}^{(1)}_{t}\mathbbm{1}\left\{\overline{X}^{(1)}_{t}\leqslant\varepsilon\right\}\right]=1$
+(4.3) $\lim_{t\downarrow 0}\frac{1}{\mu(t)}\mathbb{E}_{0}\left[\overline{X}^{(1)}_{t}\mathbbm{1}\left\{\sup_{0\leqslant s\leqslant t}|X_{s}|\leqslant\varepsilon\right\}\right]=1$
+(4.4) $\limsup_{t\downarrow 0}\frac{1}{\mu(t)}\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}|X_{s}|\mathbbm{1}\left\{\sup_{0\leqslant s\leqslant t}|X_{s}|\leqslant\varepsilon\right\}\right]\leqslant 2d$
 Proof.
 For any $\varepsilon>0$ we note that
-$\left|\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\wedge\varepsilon\right]-\mu(t)\right|\leqslant\max\{\varepsilon,1\}\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>\varepsilon\wedge 1\right).$
-Therefore, (4.1) follows as $\lim_{t\to 0}\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>\varepsilon\wedge 1\right)/\mu(t)=0$. Next, for any $t>0$, we have
-$0$ $\leqslant\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\wedge\varepsilon\right]-\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\mathbbm{1}\left\{\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\leqslant\varepsilon\right\}\right]$
-$\leqslant\varepsilon\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>\varepsilon\right),$
+$\left|\mathbb{E}_{0}\left[\overline{X}^{(1)}_{t}\wedge\varepsilon\right]-\mu(t)\right|\leqslant\max\{\varepsilon,1\}\mathbb{P}_{0}\left(\overline{X}^{(1)}_{t}>\varepsilon\wedge 1\right).$
+Therefore, (4.1) follows as $\lim_{t\downarrow 0}\mathbb{P}_{0}\left(\overline{X}^{(1)}_{t}>\varepsilon\wedge 1\right)/\mu(t)=0$. Next, for any $t>0$, we have
+$0$ $\leqslant\mathbb{E}_{0}\left[\overline{X}^{(1)}_{t}\wedge\varepsilon\right]-\mathbb{E}_{0}\left[\overline{X}^{(1)}_{t}\mathbbm{1}\left\{\overline{X}^{(1)}_{t}\leqslant\varepsilon\right\}\right]$
+$\leqslant\varepsilon\mathbb{P}_{0}\left(\overline{X}^{(1)}_{t}>\varepsilon\right),$
 and hence, (4.2) follows from (4.1) and Assumption 1.1. To prove (4.3), we observe that
-$0$ $\leqslant\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\leqslant\varepsilon\}\right]-\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}|X_{s}|\leqslant\varepsilon\}\right]$
-$=\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\leqslant\varepsilon,\sup_{0\leqslant s\leqslant t}|X_{s}|>\varepsilon\}\right]$
-$\leqslant\varepsilon\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}|X_{s}|>\varepsilon\right)\leqslant 2d\varepsilon\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>\varepsilon/\sqrt{d}\right),$
+$0$ $\leqslant\mathbb{E}_{0}\left[\overline{X}^{(1)}_{t}\mathbbm{1}\{\overline{X}^{(1)}_{t}\leqslant\varepsilon\}\right]-\mathbb{E}_{0}\left[\overline{X}^{(1)}_{t}\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}|X_{s}|\leqslant\varepsilon\}\right]$
+$=\mathbb{E}_{0}\left[\overline{X}^{(1)}_{t}\mathbbm{1}\{\overline{X}^{(1)}_{t}\leqslant\varepsilon,\sup_{0\leqslant s\leqslant t}|X_{s}|>\varepsilon\}\right]$
+$\leqslant\varepsilon\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}|X_{s}|>\varepsilon\right)\leqslant 2d\varepsilon\mathbb{P}_{0}\left(\overline{X}^{(1)}_{s}>\varepsilon/\sqrt{d}\right),$
 where the last inequality follows from the isotropy of $X_{t}$. Again by Assumption 1.1 and (4.2), we conclude that
-(4.5) $\ \ \ \ \lim_{t\to 0}\frac{1}{\mu(t)}\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}|X_{s}|\leqslant\varepsilon\}\right]$
-$=\lim_{t\to 0}\frac{1}{\mu(t)}\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\leqslant\varepsilon\}\right]=1.$
+(4.5) $\ \ \ \ \lim_{t\downarrow 0}\frac{1}{\mu(t)}\mathbb{E}_{0}\left[\overline{X}^{(1)}_{t}\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}|X_{s}|\leqslant\varepsilon\}\right]$
+$=\lim_{t\downarrow 0}\frac{1}{\mu(t)}\mathbb{E}_{0}\left[\overline{X}^{(1)}_{t}\mathbbm{1}\{\overline{X}^{(1)}_{t}\leqslant\varepsilon\}\right]=1.$
 Finally, using the symmetry of the random variables $X^{(i)}_{t}$ for $i=1,\ldots,d$, we note that
 $\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}|X^{(i)}_{s}|\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}|X^{(i)}_{s}|\leqslant\varepsilon\}\right]\leqslant\int_{0}^{\varepsilon}\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}|X^{(i)}_{s}|>u\right)du$
-$\leqslant 2\int_{0}^{\varepsilon}\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}X^{(i)}_{s}>u\right)=2\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(i)}_{s}\wedge\varepsilon\right].$
+$\leqslant 2\int_{0}^{\varepsilon}\mathbb{P}_{0}\left(\overline{X}^{(i)}_{t}>u\right)du=2\mathbb{E}_{0}\left[\overline{X}^{(i)}_{t}\wedge\varepsilon\right].$
 […]
@@
 This shows with Assumption 1.1 and (4.4) that for any $\varepsilon>0$,
-$\limsup_{t\to 0}\frac{1}{\mu(t)}\int_{S_{f}}\mathbb{E}_{x}\left[\sup_{0\leqslant s\leqslant t}|X_{s}-x|^{1+\kappa}\mathbbm{1}\left\{\sup_{0\leqslant s\leqslant t}|X_{s}-x|\leqslant 2R\right\}\right]dx\leqslant 2d\varepsilon^{\kappa}\mathrm{Vol}(S_{f}).$
+$\limsup_{t\downarrow 0}\frac{1}{\mu(t)}\int_{S_{f}}\mathbb{E}_{x}\left[\sup_{0\leqslant s\leqslant t}|X_{s}-x|^{1+\kappa}\mathbbm{1}\left\{\sup_{0\leqslant s\leqslant t}|X_{s}-x|\leqslant 2R\right\}\right]dx\leqslant 2d\varepsilon^{\kappa}\mathrm{Vol}(S_{f}).$
 Letting $\varepsilon\to 0$, the above inequality combined with (4.9) yields
-(4.10) $\lim_{t\to 0}\frac{1}{\mu(t)}\left|\mathcal{R}^{X}_{f}(t)-\int_{S_{f}}\mathbb{E}_{x}\left[\sup_{0\leqslant s\leqslant t}-\langle X_{s}-x,\nabla f(x)\rangle\mathbbm{1}\{X_{s}\in S_{f},0\leqslant s\leqslant t\}\right]dx\right|=0.$
+(4.10) $\lim_{t\downarrow 0}\frac{1}{\mu(t)}\left|\mathcal{R}^{X}_{f}(t)-\int_{S_{f}}\mathbb{E}_{x}\left[\sup_{0\leqslant s\leqslant t}-\langle X_{s}-x,\nabla f(x)\rangle\mathbbm{1}\{X_{s}\in S_{f},0\leqslant s\leqslant t\}\right]dx\right|=0.$
 Let us denote
-$\delta(x)=\inf\{|x-y|:y\in\partial S_{f}\}.$
+$\delta(x)=\inf\{|x-y|:y\notin S_{f}\}.$
 Then, any $x\in S_{f}$ and $t\geqslant 0$ we have
@@
 $\ \ \ \ \mathbb{E}_{x}\left[\sup_{0\leqslant s\leqslant t}-\langle X_{s}-x,\nabla f(x)\rangle\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}|X_{s}-x|\leqslant r\}\right]$
-$=\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}|X_{s}|\leqslant r\}\right]|\nabla f(x)|=:\mu(t,r)|\nabla f(x)|.$
+$=\mathbb{E}_{0}\left[\overline{X}^{(1)}_{t}\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}|X_{s}|\leqslant r\}\right]|\nabla f(x)|=:\mu(t,r)|\nabla f(x)|.$
 As a result, (4.11) implies the following upper and lower bound:
@@
 Combining (4.10) and (4.12), we obtain
-(4.13) $\limsup_{t\to 0}\frac{\mathcal{R}^{X}_{f}(t)}{\mu(t)}$ $\leqslant\limsup_{t\to 0}\frac{\mu(t,2R)}{\mu(t)}\int_{\Omega}|\nabla f(x)|dx$
+(4.13) $\limsup_{t\downarrow 0}\frac{\mathcal{R}^{X}_{f}(t)}{\mu(t)}$ $\leqslant\limsup_{t\downarrow 0}\frac{\mu(t,2R)}{\mu(t)}\int_{\Omega}|\nabla f(x)|dx$
 $=\int_{\mathbb{R}^{d}}|\nabla f(x)|dx,$
 where the last equality follows from (4.2) in Lemma 4.3. On the other hand, using (4.10) and (4.12) once again we get
-$\liminf_{t\to 0}\frac{\mathcal{R}^{X}_{f}(t)}{\mu(t)}\geqslant\liminf_{t\to 0}\int_{S_{f}}|\nabla f(x)|\frac{\mu(t,\delta(x))}{\mu(t)}dx.$
-Since by (4.2) in Lemma 4.3, $\frac{\mu(t,\delta(x))}{\mu(t)}\to 1$ as $t\to 0$, applying Fatou’s lemma we conclude that
-(4.14) $\liminf_{t\to 0}\frac{\mathcal{R}^{X}_{f}(t)}{\mu(t)}\geqslant\int_{S_{f}}|\nabla f(x)|dx=\int_{\mathbb{R}^{d}}|\nabla f(x)|dx.$
+$\liminf_{t\downarrow 0}\frac{\mathcal{R}^{X}_{f}(t)}{\mu(t)}\geqslant\liminf_{t\downarrow 0}\int_{S_{f}}|\nabla f(x)|\frac{\mu(t,\delta(x))}{\mu(t)}dx.$
+Since by (4.2) in Lemma 4.3, $\frac{\mu(t,\delta(x))}{\mu(t)}\to 1$ as $t\downarrow 0$, applying Fatou’s lemma we conclude that
+(4.14) $\liminf_{t\downarrow 0}\frac{\mathcal{R}^{X}_{f}(t)}{\mu(t)}\geqslant\int_{S_{f}}|\nabla f(x)|dx=\int_{\mathbb{R}^{d}}|\nabla f(x)|dx.$
 Combining (4.13) and (4.14), we conclude the proof of the theorem. ∎
@@
 Let $\Omega$ be a bounded, open set in $\mathbb{R}^{d}$ with finite perimeter. Then,
-$\liminf_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\geqslant\mathrm{Per}(\Omega).$
+$\liminf_{t\downarrow 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\geqslant\mathrm{Per}(\Omega).$
 Remark 5.2.
-When $\Omega$ has $C^{1,1}$-boundary, the above proposition can be proved using a half-space argument similar to the proof of [9, Lemma 3.1].
-However, the argument in the aforementioned paper fails when $\partial\Omega$ is not regular.
+When $\Omega$ has $C^{1,1}$ boundary, the above proposition can be proved using a half-space argument similar to the proof of [18, Lemma 3.1].
+However, the argument in the aforementioned paper requires the $R$-smooth boundary condition, which does not hold when $\partial\Omega$ is not $C^{1,1}$.
 The above result is proved using approximation argument using mollifiers.
@@
 Let us consider $f=\mathbbm{1}_{\Omega}$. By Lemma 5.3 and Theorem 4.2, for any $\varepsilon>0$,
-$\liminf_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\geqslant\liminf_{t\to 0}\frac{\mathcal{R}^{X}_{f_{\varepsilon}}(t)}{\mu(t)}=\int_{\mathbb{R}^{d}}|\nabla f_{\varepsilon}(x)|dx.$
+$\liminf_{t\downarrow 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\geqslant\liminf_{t\downarrow 0}\frac{\mathcal{R}^{X}_{f_{\varepsilon}}(t)}{\mu(t)}=\int_{\mathbb{R}^{d}}|\nabla f_{\varepsilon}(x)|dx.$
 As $\lim_{\varepsilon\to 0}\|f_{\varepsilon}-\mathbbm{1}_{\Omega}\|_{L^{1}}=0$, by lower semi-continuity of the variation it follows that
-$\liminf_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\geqslant\liminf_{\varepsilon\to 0}\int_{\mathbb{R}^{d}}|\nabla f_{\varepsilon}(x)|\mu(dx)\geqslant\|D\mathbbm{1}_{\Omega}\|=\mathrm{Per}(\Omega).$
+$\liminf_{t\downarrow 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\geqslant\liminf_{\varepsilon\to 0}\int_{\mathbb{R}^{d}}|\nabla f_{\varepsilon}(x)|dx\geqslant\|D\mathbbm{1}_{\Omega}\|=\mathrm{Per}(\Omega).$
 This completes the proof of the proposition. ∎
@@
 Then for any translation invariant isotropic process $X=(X_{t})_{t\geqslant 0}$ satisfying Assumption 1.1 we have
-$\mathrm{Per}(\Omega)\leqslant\liminf_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\leqslant\limsup_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\leqslant\frac{\sup_{\partial\Omega}|\nabla\phi|}{\inf_{\partial\Omega}|\nabla\phi|}\mathrm{Per}(\Omega).$
+$\mathrm{Per}(\Omega)\leqslant\liminf_{t\downarrow 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\leqslant\limsup_{t\downarrow 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\leqslant\frac{\sup_{\partial\Omega}|\nabla\phi|}{\inf_{\partial\Omega}|\nabla\phi|}\mathrm{Per}(\Omega).$
 The theorem is proved in several intermediate steps. First, we note the following result about continuity of the perimeters of level sets.
@@
 For any $\varepsilon>0$,
-$\lim_{t\to 0}\frac{1}{\mu(t)}\int_{\Omega_{\varepsilon}}\mathbb{P}_{x}\left(\inf_{0\leqslant s\leqslant t}\phi(X_{s})\leqslant 0\right)dx=0$
+$\lim_{t\downarrow 0}\frac{1}{\mu(t)}\int_{\Omega_{\varepsilon}}\mathbb{P}_{x}\left(\inf_{0\leqslant s\leqslant t}\phi(X_{s})\leqslant 0\right)dx=0$
 Proof.
@@
 where $M=\|\nabla\phi\|_{\infty}$ and $L^{\prime}=L/(1+\kappa)$. As a result, for any $x\in\Omega_{\varepsilon}$, Assumption 1.1 implies that
-$\ \ \ \ \lim_{t\to 0}\frac{1}{\mu(t)}\int_{\Omega_{\varepsilon}}\mathbb{P}_{x}\left(\inf_{0\leqslant s\leqslant t}\phi(X_{s})\leqslant 0\right)dx$
-$\leqslant\lim_{t\to 0}\frac{1}{\mu(t)}\left(\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}|X_{s}|>\frac{\varepsilon}{2M}\right)+\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}|X_{s}|>\left(\frac{\varepsilon}{L^{\prime}}\right)^{1/1+\kappa}\right)\right)\mathrm{Vol}(\Omega)$
+$\ \ \ \ \lim_{t\downarrow 0}\frac{1}{\mu(t)}\int_{\Omega_{\varepsilon}}\mathbb{P}_{x}\left(\inf_{0\leqslant s\leqslant t}\phi(X_{s})\leqslant 0\right)dx$
+$\leqslant\lim_{t\downarrow 0}\frac{1}{\mu(t)}\left(\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}|X_{s}|>\frac{\varepsilon}{2M}\right)+\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}|X_{s}|>\left(\frac{\varepsilon}{L^{\prime}}\right)^{1/1+\kappa}\right)\right)\mathrm{Vol}(\Omega)$
 $=0.$
@@
 For any $\eta\in(0,1)$ and $\varepsilon>0$,
-$\limsup_{t\to 0}\frac{1}{\mu(t)}\int_{\Omega^{\prime}_{\varepsilon}}\mathbb{P}_{x}\left(\inf_{0\leqslant s\leqslant t}\phi(X_{s})\leqslant 0\right)dx$
+$\limsup_{t\downarrow 0}\frac{1}{\mu(t)}\int_{\Omega^{\prime}_{\varepsilon}}\mathbb{P}_{x}\left(\inf_{0\leqslant s\leqslant t}\phi(X_{s})\leqslant 0\right)dx$
 $\leqslant\sup_{0\leqslant v\leqslant\varepsilon}\mathcal{H}^{d-1}(\partial\Omega_{v})\frac{M_{\varepsilon}}{\eta m_{\varepsilon}}+2\sup_{0\leqslant v\leqslant\varepsilon}\mathcal{H}^{d-1}(\partial\Omega_{v})\frac{(1-\eta)^{-\frac{\kappa}{\kappa+1}}}{m_{\varepsilon}}\varepsilon^{\frac{\kappa}{\kappa+1}},$
@@
 (6.4) $\ \ \ \ \int_{\Omega^{\prime}_{\varepsilon}}\mathbb{P}_{x}\left(\sup_{0\leqslant s\leqslant t}\langle X_{s}-x,\nabla\phi(x)\rangle>\eta\phi(x)\right)dx$
-$=\int_{0}^{\varepsilon}\int_{\partial\Omega_{v}}\mathbb{P}_{0}\left(|\nabla\phi(x)|\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>\eta v\right)\frac{\mathcal{H}^{d-1}(dx)}{|\nabla\phi(x)|}dv.$
+$=\int_{0}^{\varepsilon}\int_{\partial\Omega_{v}}\mathbb{P}_{0}\left(|\nabla\phi(x)|\overline{X}^{(1)}_{t}>\eta v\right)\frac{\mathcal{H}^{d-1}(dx)}{|\nabla\phi(x)|}dv.$
 Then, (6.4) yields
 $\ \ \ \ \int_{\Omega^{\prime}_{\varepsilon}}\mathbb{P}_{x}\left(\sup_{0\leqslant s\leqslant t}\langle X_{s}-x,\nabla\phi(x)\rangle>\eta\phi(x)\right)dx$
-$\leqslant\sup_{0\leqslant v\leqslant\varepsilon}\mathcal{H}^{d-1}(\partial\Omega_{v})\frac{1}{m_{\varepsilon}}\int_{0}^{\varepsilon}\mathbb{P}_{0}\left(M_{\varepsilon}\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>\eta v\right)dv$
-$=\sup_{0\leqslant v\leqslant\varepsilon}\mathcal{H}^{d-1}(\partial\Omega_{v})\frac{M_{\varepsilon}}{\eta m_{\varepsilon}}\mathbb{E}_{0}\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\wedge\frac{\eta\varepsilon}{M_{\varepsilon}}\right)$
+$\leqslant\sup_{0\leqslant v\leqslant\varepsilon}\mathcal{H}^{d-1}(\partial\Omega_{v})\frac{1}{m_{\varepsilon}}\int_{0}^{\varepsilon}\mathbb{P}_{0}\left(M_{\varepsilon}\overline{X}^{(1)}_{t}>\eta v\right)dv$
+$=\sup_{0\leqslant v\leqslant\varepsilon}\mathcal{H}^{d-1}(\partial\Omega_{v})\frac{M_{\varepsilon}}{\eta m_{\varepsilon}}\mathbb{E}_{0}\left(\overline{X}^{(1)}_{t}\wedge\frac{\eta\varepsilon}{M_{\varepsilon}}\right)$
 By (4.1) in Lemma 4.3, we get
-(6.5) $\limsup_{t\to 0}\frac{1}{\mu(t)}\int_{\Omega^{\prime}_{\varepsilon}}\mathbb{P}_{x}\left(\sup_{0\leqslant s\leqslant t}\langle X_{s}-x,\nabla\phi(x)\rangle>\eta\phi(x)\right)dx$
+(6.5) $\limsup_{t\downarrow 0}\frac{1}{\mu(t)}\int_{\Omega^{\prime}_{\varepsilon}}\mathbb{P}_{x}\left(\sup_{0\leqslant s\leqslant t}\langle X_{s}-x,\nabla\phi(x)\rangle>\eta\phi(x)\right)dx$
 $\leqslant\sup_{0\leqslant v\leqslant\varepsilon}\mathcal{H}^{d-1}(\partial\Omega_{v})\frac{M_{\varepsilon}}{\eta m_{\varepsilon}}$
@@
 $\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}|X_{s}|^{1+\kappa}\wedge(1-\eta)\varepsilon\right]=\mathbb{E}_{0}\left[\left(\sup_{0\leqslant s\leqslant t}|X_{s}|\wedge(1-\eta)^{\frac{1}{1+\kappa}}\varepsilon^{\frac{1}{1+\kappa}}\right)^{\kappa+1}\right]$
-$\leqslant((1-\eta)\varepsilon)^{\frac{1}{\kappa+1}}\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}|X_{s}|\wedge(1-\eta)^{\frac{1}{1+\kappa}}\varepsilon^{\frac{1}{1+\kappa}}\right]\leqslant 2\mu(t)((1-\eta)\varepsilon)^{\frac{\kappa}{\kappa+1}}$
+$\leqslant((1-\eta)\varepsilon)^{\frac{\kappa}{\kappa+1}}\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}|X_{s}|\wedge(1-\eta)^{\frac{1}{1+\kappa}}\varepsilon^{\frac{1}{1+\kappa}}\right]\leqslant 2\mu(t)((1-\eta)\varepsilon)^{\frac{\kappa}{\kappa+1}}$
 for any $\varepsilon\in(0,1)$. Since $m_{\varepsilon}$ is bounded away from 0 as $\varepsilon\to 0$, (6.6) implies
-(6.7) $\limsup_{t\to 0}\frac{1}{\mu(t)}\int_{\Omega^{\prime}_{\varepsilon}}\mathbb{P}_{x}\left(\sup_{0\leqslant s\leqslant t}|X_{s}-x|^{1+\kappa}>(1-\eta)\phi(x)\right)dx$
+(6.7) $\limsup_{t\downarrow 0}\frac{1}{\mu(t)}\int_{\Omega^{\prime}_{\varepsilon}}\mathbb{P}_{x}\left(\sup_{0\leqslant s\leqslant t}|X_{s}-x|^{1+\kappa}>(1-\eta)\phi(x)\right)dx$
 $\leqslant 2\sup_{0\leqslant v\leqslant\varepsilon}\mathcal{H}^{d-1}(\partial\Omega_{v})\frac{(1-\eta)^{-\frac{\kappa}{\kappa+1}}}{m_{\varepsilon}}\varepsilon^{\frac{\kappa}{\kappa+1}}.$
@@
 By Proposition 5.1, it suffices to prove that
-$\limsup_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\leqslant\frac{\sup_{\partial\Omega}|\nabla\phi|}{\inf_{\partial\Omega}|\nabla\phi|}\mathrm{Per}(\Omega).$
+$\limsup_{t\downarrow 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\leqslant\frac{\sup_{\partial\Omega}|\nabla\phi|}{\inf_{\partial\Omega}|\nabla\phi|}\mathrm{Per}(\Omega).$
 For any $\varepsilon>0$ we have
-(6.8) $\ \ \ \ \limsup_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}=\int_{\Omega}\mathbb{P}_{x}(\tau^{X}_{\Omega}\leqslant t)dx$
-$\leqslant\limsup_{t\to 0}\frac{1}{\mu(t)}\int_{\Omega_{\varepsilon}}\mathbb{P}_{x}\left(\inf_{0\leqslant s\leqslant t}\phi(X_{s})\leqslant 0\right)dx$
-$+\limsup_{t\to 0}\frac{1}{\mu(t)}\int_{\Omega^{\prime}_{\varepsilon}}\mathbb{P}_{x}\left(\inf_{0\leqslant s\leqslant t}\phi(X_{s})\leqslant 0\right)dx.$
+(6.8) $\ \ \ \ \limsup_{t\downarrow 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}=\int_{\Omega}\mathbb{P}_{x}(\tau^{X}_{\Omega}\leqslant t)dx$
+$\leqslant\limsup_{t\downarrow 0}\frac{1}{\mu(t)}\int_{\Omega_{\varepsilon}}\mathbb{P}_{x}\left(\inf_{0\leqslant s\leqslant t}\phi(X_{s})\leqslant 0\right)dx$
+$+\limsup_{t\downarrow 0}\frac{1}{\mu(t)}\int_{\Omega^{\prime}_{\varepsilon}}\mathbb{P}_{x}\left(\inf_{0\leqslant s\leqslant t}\phi(X_{s})\leqslant 0\right)dx.$
 By Lemma 6.3, the first term on the right hand side of (6.8) is zero. Recalling Lemma 6.4, we get
-$\ \ \ \ \limsup_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}$
+$\ \ \ \ \limsup_{t\downarrow 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}$
 $\leqslant 2\sup_{0\leqslant v\leqslant\varepsilon}\mathcal{H}^{d-1}(\partial\Omega_{v})\frac{M_{\varepsilon}}{\eta m_{\varepsilon}}+2\sup_{0\leqslant v\leqslant\varepsilon}\mathcal{H}^{d-1}(\partial\Omega_{v})\frac{(1-\eta)^{-\frac{\kappa}{\kappa+1}}}{m_{\varepsilon}}\varepsilon^{\frac{\kappa}{\kappa+1}}$
@@
 Finally, letting $\eta\to 1$, we conclude that
-$\limsup_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\leqslant\frac{\sup_{\partial\Omega}|\nabla\phi|}{\inf_{\partial\Omega}|\nabla\phi|}\mathcal{H}^{d-1}(\partial\Omega)=\frac{\sup_{\partial\Omega}|\nabla\phi|}{\inf_{\partial\Omega}|\nabla\phi|}\mathrm{Per}(\Omega).$
+$\limsup_{t\downarrow 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\leqslant\frac{\sup_{\partial\Omega}|\nabla\phi|}{\inf_{\partial\Omega}|\nabla\phi|}\mathcal{H}^{d-1}(\partial\Omega)=\frac{\sup_{\partial\Omega}|\nabla\phi|}{\inf_{\partial\Omega}|\nabla\phi|}\mathrm{Per}(\Omega).$
 This completes the proof of the theorem. ∎
 Concluding proof of Theorem 1.1
-The final step of the proof of Theorem 1.1 is to construct a function $\phi$ satisfying conditions of Theorem 6.1 for domains with $C^{1,1}$-boundary.
+The final step of the proof of Theorem 1.1 is to construct a function $\phi$ satisfying conditions of Theorem 6.1 for domains with $C^{1,1}$ boundary.
 Lemma 6.5.
-Let $\Omega\subset\mathbb{R}^{d}$ be a bounded open set with $C^{1,1}$-boundary. Then there exists a bounded $C^{1,1}$ function $\phi:\mathbb{R}^{d}\to\mathbb{R}$ such that
+Let $\Omega\subset\mathbb{R}^{d}$ be a bounded open set with $C^{1,1}$ boundary. Then there exists a bounded $C^{1,1}$ function $\phi:\mathbb{R}^{d}\to\mathbb{R}$ such that
   1. (1)
@@
 $\delta(x)=\begin{cases}d(x,\partial\Omega)&\mbox{if $x\in\Omega$}\\ -d(x,\partial\Omega)&\mbox{if $x\in\Omega^{c}$}.\end{cases}$
-Since $\Omega$ has $C^{1,1}$-boundary, there exists $r>0$ such that $\delta$ is $C^{1,1}$ in $\Omega^{\prime}_{r}:=\{x\in\mathbb{R}^{d}:|\delta(x)|<r\}$ and $|\nabla\delta|=1$ in $\Omega^{\prime}_{r}$.
-Define a $C^{1,1}$ function $\varphi:\mathbb{R}\to\mathbb{R}$ such that $\varphi(x)=x$ when $|x|<r/2$ and equals $r$ when $|x|>r$ and consider $\phi=\varphi\circ\delta$.
-Clearly, $\phi$ is $C^{1,1}$ and $\phi(x)>0$ if and only if $x\in\Omega$.
-Also, for any $x\in\Omega^{\prime}_{r/2}$, $|\nabla\phi(x)|=|\varphi^{\prime}(\delta(x))\nabla\delta(x)|=1$.
+Since $\Omega$ has $C^{1,1}$ boundary, there exists $r_{0}>0$ such that $\delta$ is $C^{1,1}$ in $\Omega^{\prime}_{r_{0}}:=\{x\in\mathbb{R}^{d}:|\delta(x)|<r_{0}\}$ and $|\nabla\delta|=1$ in $\Omega^{\prime}_{r_{0}}$.
+Fix $0<r<r_{0}$ and define a $C^{1,1}$ function $\varphi:\mathbb{R}\to\mathbb{R}$
+$\varphi(x)=\begin{cases}x&\mbox{if $|x|<r/2$}\\ r&\mbox{if $x\geqslant r$}\\ -r&\mbox{if $x\leqslant-r$}.\end{cases}$
+Consider the function $\phi=\varphi\circ\delta$.
+As $\delta$ is $C^{1,1}$ in $\Omega^{\prime}_{r_{0}}$ and $\phi$ is constant on $\{x:\delta(x)\geqslant r\}$, $\phi$ is a bounded $C^{1,1}$ function on $\mathbb{R}^{d}$.
+Also, $\phi(x)>0$ if and only if $x\in\Omega$.
+For any $x\in\Omega^{\prime}_{r/2}$, $|\nabla\phi(x)|=|\varphi^{\prime}(\delta(x))\nabla\delta(x)|=1$.
 This completes the proof of the lemma.
@@
 Remark 6.6.
-The $C^{1,1}$-regularity assumption on $\partial\Omega$ cannot be weakened.
-From [8, Corollary 6] it is known that the differentiability of $\delta$ in a neighborhood of $\partial\Omega$ implies that $\partial\Omega$ is $C^{1,1}$.
+The $C^{1,1}$ regularity assumption on $\partial\Omega$ cannot be weakened.
+From [17, Corollary 6] it is known that the differentiability of $\delta$ in a neighborhood of $\partial\Omega$ implies that $\partial\Omega$ is $C^{1,1}$.
 We note that $\phi$ in Lemma 6.5 satisfies the conditions of Theorem 6.1 with $\kappa=1$, and $\Omega=\phi^{-1}(0,\infty)$.
@@
 ∎
-Proof of Corollary 1.3.
-For any $\varepsilon>0$, using Markov inequality we obtain
-$\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>\varepsilon\right)\leqslant\frac{\mathbb{E}_{0}\left[\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\right)^{\alpha}\right]}{\varepsilon^{\alpha}}.$
-Also, from the given conditions and noting that $\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\geqslant 0$ if $X_{0}=0$, we have
-$0\leqslant 1-\frac{\mu(t)}{m(t)}\leqslant\frac{\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\mathbbm{1}\left\{\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>1\right\}\right]}{m(t)}\leqslant\frac{\mathbb{E}_{0}\left[\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\right)^{\alpha}\right]}{m(t)},$
-and the right hand side goes to 0 as $t\to 0$. Hence, $\mu(t)/m(t)\to 1$ as $t\to 0$ and Assumption 1.1 is satisfied. The proof is concluded by applying Theorem 1.1. ∎
+Proof of Proposition 1.3.
+Since $\mathbb{P}_{0}(\overline{X}^{(1)}_{t}\geqslant 0)=1$, for any $\varepsilon>0$, using Markov inequality we obtain
+$\mathbb{P}_{0}\left(\overline{X}^{(1)}_{t}>\varepsilon\right)\leqslant\frac{\mathbb{E}_{0}\left[\left(\overline{X}^{(1)}_{t}\right)^{p}\right]}{\varepsilon^{p}}.$
+Also, from the given conditions we have
+$0\leqslant 1-\frac{\mu(t)}{m(t)}\leqslant\frac{\mathbb{E}_{0}\left[\overline{X}^{(1)}_{t}\mathbbm{1}\left\{\overline{X}^{(1)}_{t}>1\right\}\right]}{m(t)}\leqslant\frac{\mathbb{E}_{0}\left[\left(\overline{X}^{(1)}_{t}\right)^{p}\right]}{m(t)},$
+and the right hand side goes to 0 as $t\downarrow 0$, where $m(t)=\mathbb{E}_{0}[\overline{X}^{(1)}_{t}$].
+Hence, $\mu(t)/m(t)\to 1$ as $t\downarrow 0$ and Assumption 1.1 is satisfied.
+The proof is concluded by applying Theorem 1.1.
+∎